### v1

# Disturbance Compensation for Safe Kinematic Control of Robotic Systems with Closed Architecture

Abstract
In commercial robotic systems, it is common to encounter a closed inner-loop torque controller that is not user-modifiable. However, the outer-loop controller, which sends kinematic commands such as position or velocity for the inner-loop controller to track, is typically exposed to users. In this work, we focus on the development of an easily integrated add-on at the outer-loop layer by combining disturbance rejection control and robust control barrier function for high-performance tracking and safe control of the whole dynamic system of an industrial manipulator. This is particularly beneficial when 1) the inner-loop controller is imperfect, unmodifiable, and uncertain; and 2) the dynamic model exhibits significant uncertainty. Stability analysis, formal safety guarantee proof, and hardware experiments with a PUMA robotic manipulator are presented. Our solution demonstrates superior performance in terms of simplicity of implementation, robustness, tracking precision, and safety compared to the state of the art. Video: https://youtu.be/zw1tanvrV8Q
I Introduction
Robotic systems typically use a hierarchical architecture combining perception, planning, and low-level control, which improves modularity and reliability. In the control layer, a kinematic controller (outer loop) and a dynamic controller (inner loop) are commonly used. Since tuning the inner loop requires expertise, it is usually hidden from users for safety, forming a closed architecture [1, 2, 3, 4]. Users can thus only design the kinematic controller, sending position or velocity commands. However, wear, modifications, or payload changes can alter dynamics and degrade inner-loop performance [5]. This work addresses the challenge of achieving unified high-performance and safe tracking control using only the kinematic controller, despite an imperfect, unmodifiable inner loop, and uncertain dynamics.
Fig. 1 shows that our approach successfully follows an infinity-shaped reference trajectory despite various uncertainties and a poor inner-loop unknown controller (see the blue trajectory). In contrast, using the poor inner-loop controller without our approach’s compensation results in poor trajectory tracking (see the red trajectory).
Figure 1: Trajectory tracking of the PUMA 500 for an infinity-shaped reference.
A detailed literature review can be found in the subsequent section. Existing works typically rely on the following assumptions: 1) The inner-loop controller is known and accessible to the user; 2) The inner-loop controller is known but not accessible to the user; 3) The inner-loop controller has perfect tracking for outer-loop commands. Our goal is to minimize reliance on knowledge of the dynamic model and the inner-loop controller, while still guaranteeing tracking and safe control using only an outer-loop kinematic controller.
As shown in Fig. 2, existing robotic systems typically send desired position ($q^{*}$) or velocity ($\dot{q}^{*}$) commands to the inner-loop controller. Our solution, inserted between them (gray box), comprises a disturbance-rejection tracking controller, an extended state observer (ESO), and an extended state observer-based robust quadratic programming (ESOR-QP) module. The ESO estimates the total disturbance $\hat{f}$ using the joint angle $q$ and the safe kinematic control $q_{d,\text{safe}}$ or $\dot{q}_{d,\text{safe}}$. This estimate is compensated in the tracking controller to produce nominal commands $q_{d}$ or $\dot{q}_{d}$, which are monitored by the ESOR-QP to enforce robust safety constraints.
The contributions of our work are summarized as follows:
  • We innovatively formulate the kinematic control of closed-architecture robotic manipulator systems as a disturbance rejection problem, requiring only a practical assumption on the structure, Jacobian, and nominal inertia matrix.
  • We develop a robust control barrier function (CBF) framework based on an extended state observer, enabling safe control of general nonlinear affine systems with high relative degrees under disturbances. To the best of our knowledge, this is the first work achieving safe kinematic control of closed-architecture robots under largely unknown uncertainties.
  • As shown in Fig. 2, our method is implemented as a standalone add-on module that enhances existing robotic manipulators without modifying the inner-loop controller. Real-time control at 1 kHz is demonstrated on a PUMA 500 manipulator.
Figure 2: The proposed framework in this paper, is marked as the Outer-loop Controller. The total estimated disturbance $\hat{f}$ is used for both disturbance compensation and safe control using ESOR-QP.
The rest of the paper is organized as follows: Section II reviews the related work. Section III introduces the preliminaries and problem formulation. Section IV presents the proposed framework. Hardware experiments are reported in Section V. Concluding remarks are in Section VI.
II Related Work
An overview of the comparison between our work and the most related works is presented in Table I.
II-A Adaptive Control for Closed Architecture
Learning-based Adaptive Control: [2, 3, 4] employ neural network–based outer-loop adaptation to approximate the kinematics, dynamics, and inner-loop controller. As summarized in Table I, they make the weakest assumption about the inner loop, identical to ours, where $\Psi(\cdot)$ denotes an unknown structure possibly containing proportional, integral, or state-dependent terms. However, they rely on numerous tuning parameters and limited interpretability.
Conventional Adaptive Control: [1] introduces dynamic modularity with an adaptive controller under a moderate assumption of a known PI/PID inner-loop structure.
These studies address tracking control only, without safety considerations. In our experiments, we compare tracking performance with the learning-based adaptive control due to the shared inner-loop assumption.
TABLE I: Overview of our work’s position in the literature
Methods Tracking Safety
Assumption of
Inner-loop Controller
Ref.
Learning-based
Adaptive Control
Weak
$\tau=-K_{d}(\dot{q}-\dot{q}_{d})+\Psi(q,\dot{q},q_{d},\int_{0}^{t}q\,dt,\int_{0}^{t}q_{d}\,dt)$
[2, 3, 4]
Conventional
Adaptive Control
Moderate, known PID/PI structure
[1]
Reduced-order CBF -
Strong, known PD structure,
perfect tracking or
bounded error
[6, 7, 8]
Robust CBF - - [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]
Ours
Weak
$\tau=-K_{d}(\dot{q}-\dot{q}_{d})+\Psi(q,\dot{q},q_{d},\int_{0}^{t}q\,dt,\int_{0}^{t}q_{d}\,dt)$
  • The two dashes for CBFs’ tracking indicate that we only compare our safety performance with them, as their nominal tracking controllers vary significantly across papers. The third dash for the robust CBF indicates that those works do not consider two-loop controllers.
II-B Reduced-order Model-based CBF
Many CBF-based safety approaches rely on high-fidelity dynamic models with torque control, yet obtaining such models remains challenging. Reduced-order or kinematic CBFs have been introduced to ensure safety with lower-fidelity models. For instance, [6] guarantees safety via safe velocity control but assumes a perfect inner-loop controller capable of instant tracking. Similarly, [7, 8] employ kinetic energy–based CBFs that offer partial robustness to model uncertainty but still depend on numerous dynamic parameters and a known PD-type inner loop. In contrast, our work addresses closed-architecture robots with minimal knowledge of the inner-loop controller and significant model uncertainty.
II-C Robust CBF
Robust CBFs have recently been developed to handle model uncertainties and external disturbances [9, 10, 11, 12, 13, 14, 15, 17, 19, 20, 21, 25, 18, 16, 23, 24, 22]. Among them, disturbance observer–based robust CBFs (DOB-CBFs) [25, 18, 16, 23, 24, 22, 15, 17, 19, 20, 21] enable active disturbance estimation without requiring worst-case bounds. Unlike these methods where estimation error bounds are absent or loosely derived via Lyapunov analysis, our approach obtains them directly from the observer error dynamics, resulting in tighter and less conservative bounds. Experimental comparisons with state-of-the-art methods (e.g., [16]) further demonstrate this advantage. In contrast to our conference version [25] that only considered relative-degree-one systems, this work provides a generalized framework with rigorous proofs guaranteeing the forward invariance of the safety set.
III Preliminary and Problem Formulation
III-A System Dynamics
We consider a general robotic system with an Euler–Lagrange formulation incorporating a generalized external disturbance force $F_{ext}$:
$\begin{cases}{M}\ddot{q}+{C}\dot{q}+{G}+F_{r}=\tau+{J}^{T}F_{ext}\\ \hfil y=\zeta(q)\end{cases},$ (1)
where $q,\dot{q},\ddot{q}\in\mathbb{R}^{n}$ are the joint position, velocity, and acceleration, respectively. ${M}\in\mathbb{R}^{n\times n}$ is the symmetric positive-definite inertia matrix; ${C}\in\mathbb{R}^{n\times n}$ denotes centrifugal and Coriolis term; ${G}\in\mathbb{R}^{n}$ is the gravity term; $F_{r}\in\mathbb{R}^{n}$ accounts for the friction torque; $\tau\in\mathbb{R}^{n}$ represents the torque (provided by the inner-loop controller); ${J}\in\mathbb{R}^{6\times n}$ is the Jacobian matrix; and $F_{ext}\in\mathbb{R}^{6}$ represents a generalized external disturbance force. The output $y\in\mathbb{R}^{m}$, and $\zeta:\mathbb{R}^{n}\to\ \mathbb{R}^{m}$ is the mapping from joint space to task space.
The most common inner-loop controller for robot manipulators typically consists of a PID controller or PD controller, often combined with dynamic compensation, such as gravity. For stability, most inner-loop controllers include a derivative term for damping, $(\dot{q}-\dot{q}_{d})$, where $\dot{q}_{d}$ is a joint velocity command.
Assumption 1
We assume the inner-loop controller has the following structure:
$\tau=-K_{d}(\dot{q}-\dot{q}_{d})+\Psi(q,\dot{q},q_{d},\int_{0}^{t}q\,dt,\int_{0}^{t}q_{d}\,dt),$ (2)
with $K_{d}$ a proportional control gain, and $\Psi(\cdot)$ is a general, unknown function that may include proportional, integral, derivative, and dynamic compensation terms. To the best of our knowledge, this assumption is the most general in the literature (see Table I), requiring minimal knowledge of the inner-loop controller, similar to [2], and more general than [1].
Assumption 2
The kinematics of the system are assumed to be known for control in task space.
The system dynamic model combining (1) and (2) is:
$\begin{split}\ddot{q}&={M}^{-1}\Bigl(-{C}\dot{q}-{G}-F_{r}+{J}^{T}F_{ext}-{K}_{d}\dot{q}+\Psi(\cdot)+{K}_{d}\dot{q}_{d}\Bigr)\end{split}.$ (3)
We have removed torque control, as it is not accessible in a closed architecture. $\dot{q}_{d}$ is the kinematic control to be manipulated in the outer-loop layer, and its control law will be designed. Another reason for choosing $\dot{q}_{d}$ instead of $q_{d}$ as the control input is to avoid differentiation-induced noise.
III-B Robot Kinematic Control with Known Model
Neglecting $F_{r}$ and $F_{ext}$ in (1), the exact system dynamics is used to design the control law:
$\begin{split}\dot{q}_{d}&={K}_{d}^{-1}\Bigl({M}u_{0}+{C}\dot{q}+{G}+{K}_{d}\dot{q}-\Psi(\cdot)\Bigr)\end{split},$ (4)
where the position control input $u_{0}$ is defined as follows:
$\begin{split}u_{0}&=\ddot{q}^{\star}+k_{p}(q^{\star}-q)+k_{d}(\dot{q}^{\star}-\dot{q})\\ \end{split},$ (5)
where $\ddot{q}^{\star},\dot{q}^{\star},q^{\star}\in\mathbb{R}^{n}$ represent the desired acceleration, velocity, and position of a reference joint trajectory generated by an upstream planner. $k_{p}\in\mathbb{R}^{n\times n}$ and $k_{d}\in\mathbb{R}^{n\times n}$ are the proportional and derivative gain matrices, respectively. $\ddot{q}^{\star}$ serves only as a feedforward term, while $u_{0}$ is a position controller. By substituting (4) and (5) into (3), we get:
$\ddot{q}=\ddot{q}^{\star}+k_{p}(q^{\star}-q)+k_{d}(\dot{q}^{\star}-\dot{q}).$ (6)
Proper $k_{p}$ and $k_{d}$ make q{q} converge to ${q}^{\star}$ asymptotically.
III-C Control Barrier Functions
The system (3) can be rewritten as a nonlinear control-affine system when ignoring the disturbance vector $F_{ext}$:
$\dot{x}=\psi(x)+g(x)u,$ (7)
where $x\triangleq[q,\dot{q}]^{T}\in\mathcal{X}\subset\mathbb{R}^{2n}$, $\psi:\mathbb{R}^{2n}\rightarrow\mathbb{R}^{2n}$ and $g:\mathbb{R}^{2n}\rightarrow\mathbb{R}^{2n\times n}$ are Lipschitz continuous, and $u\triangleq\dot{q}_{d}\in\mathcal{U}\subset\mathbb{R}^{n}$ is a control input vector.
The safety of system (7) can be guaranteed using a safety set. A set $\mathcal{C}$ is considered to be a safety set if it is forward invariant in the state space $\mathcal{X}$, i.e., for system (7) if solutions for some $u\in\mathcal{U}$ starting at any initial safe condition $x(0)\in\mathcal{C}$ satisfy $x(t)\in\mathcal{C}$, $\forall t\geq 0$. The safety set $\mathcal{C}$ is defined as a 0-superlevel set of a continuously-differential function $h(x):\mathbb{R}^{2n}\rightarrow\mathbb{R}$ as:
$\mathcal{C}=\{x\in\mathbb{R}^{2n}~|~h(x)\geq 0\}.$ (8)
The function $h$ is used to synthesize a controller with safety guarantees via a CBF.
Definition 1
(Exponential CBF (ECBF) [26, 27]) Consider system (7) with relative degree $r$ for an $r$-times continuously differentiable function $h$, i.e., $L_{g}L_{\psi}h(x)=\cdots=L_{g}L_{\psi}^{r-2}h(x)=0$ and $L_{g}L_{\psi}^{r-1}h(x)\neq 0,\forall x\in\mathcal{C}$. $h(x)$ is an ECBF if there exists a row vector $K_{a}\in\mathbb{R}^{r}$ satisfying $\forall x\in\mathcal{C}$
$\sup\limits_{u\in\mathcal{U}}\big(L_{\psi}^{r}h(x)+L_{g}L_{\psi}^{r-1}h(x)u\big)\geq-K_{a}\eta_{b}(x),$ (9)
where $\eta_{b}(x)=[h(x),\dot{h}(x),\cdots,h^{(r-1)}(x)]^{T}$, $K_{a}=[k_{1},\cdots,k_{r}]$, and the values of $k_{1},\cdots,k_{r}$ satisfy specific properties given in [26, 27].
To guarantee safety, the control problem is formulated as a quadratic program (QP) with a CBF as a hard constraint [28]:
$\begin{split}u^{*}(x)&=\operatorname*{arg\,min}_{u\in U}\|u-k(x)\|^{2}\\ \text{s.t.}\quad L_{\psi}^{r}h(x)&+L_{g}L_{\psi}^{r-1}h(x)u\geq-K_{a}\eta_{b}(x),\end{split}$ (10)
where $k(x)$ is a nominal control law.
Research Objective: Given the system in (3), the nominal dynamics with limited knowledge of the system and inner-loop controller are defined as
$\ddot{q}=\bar{M}^{-1}\!\left(-\bar{C}\dot{q}-\bar{G}+\bar{K}_{d}\dot{q}_{d}\right),$ (11)
where $\bar{M}$, $\bar{C}$, and $\bar{G}$ are nominal parameters, and $\bar{K}_{d}$ is the nominal inner-loop gain. The goal is to design a $\dot{q}_{d}$ controller that achieves accurate trajectory tracking and task-space safety (e.g., collision avoidance) under an unknown inner-loop controller, uncertain dynamics, and external disturbances.
Remark 1
$\bar{C}$ and $\bar{G}$ can be a zero matrix and a zero vector, respectively, with minimal knowledge of the system. We use $\bar{C}$ and $\bar{G}$ here to make the formulation more general. The model (11) can be further reduced to:
$\ddot{q}=\bar{M}^{-1}\bar{K}_{d}\dot{q}_{d}.$ (12)
We use the nominal model (11) for general mathematical derivation, whereas the reduced nominal model (12) is employed in hardware experiments to evaluate the tracking performance and safety under limited information.
IV Proposed Framework
IV-A Robot Kinematic Control with Nominal Model
The control law in (4) is ideal, assuming known dynamics and an inner-loop controller. To handle uncertainties, we propose the following practical law based on nominal dynamics:
$\begin{split}\dot{q}_{d}&=\bar{K}_{d}^{-1}\Bigl(\bar{M}u_{0}+\bar{C}\dot{q}+\bar{G}\Bigr).\end{split}$ (13)
The following part shows the dynamics of the close-loop system with our nominal control law (13). We define the parametric discrepancies as:
$\begin{split}\Delta K_{d}&\triangleq K_{d}\bar{K}_{d}^{-1},\\ \Delta{M}&\triangleq\Delta K_{d}\bar{M}-{M},\\ \Delta{C}&\triangleq\Delta K_{d}\bar{C}-{C},\\ \Delta{G}&\triangleq\Delta K_{d}\bar{G}-{G}.\end{split}$ (14)
By rearranging (3), we get:
[…]
Substituting (13) and (14) into (15), the right-hand side of (15) becomes:
[…]
The left-hand side of (15) becomes:
$\begin{split}&\Bigl(\Delta K_{d}\bar{M}-\Delta{M}\Bigr)\ddot{q}+\Bigl(\Delta K_{d}\bar{C}-\Delta{C}\Bigr)\dot{q}\\ &+\Bigl(\Delta K_{d}\bar{G}-\Delta{G}\Bigr)+F_{r}.\end{split}$ (17)
The resulting close-loop system is:
$\begin{split}\ddot{q}=\ddot{q}^{\star}+k_{p}(q^{\star}-q)+k_{d}(\dot{q}^{\star}-\dot{q})+f.\end{split}$ (18)
where
$\begin{split}f=(\Delta K_{d}\bar{M})^{-1}\Bigl(&\Delta{M}\ddot{q}+\Delta{C}\dot{q}+\Delta{G}-F_{r}\\ &+\Psi(\cdot)+{J}^{T}F_{ext}-K_{d}\dot{q}\Bigr).\end{split}$ (19)
Let $f$ denote the total disturbance, including internal uncertainties and external disturbances. Compared with (6), $f$ degrades system performance. To counter this, an ESO [29] is employed to estimate $f$ for disturbance rejection.
IV-B Extended State Observer Design
For the $i$-th joint of the robotic manipulator (11), we have the following second-order subsystem:
$\begin{cases}\begin{array}[]{l}\dot{x}_{1i}=x_{2i}\\ \dot{x}_{2i}=F_{i}(x)+G_{i}(x)\dot{q}_{di}+f_{i},\end{array}\end{cases}$ (20)
where $x_{1i}=q_{i}$; $x_{2i}=\dot{q}_{i}$; $F_{i}(x)$, $G_{i}(x)$, and $f_{i}$ are the $i$-th row of $F=\bar{M}^{-1}(-\bar{C}\dot{q}-\bar{G})\in\mathbb{R}^{n}$, $G=\bar{M}^{-1}\bar{K}_{d}\in\mathbb{R}^{n}$, and $f$, respectively.
By treating $f_{i}$ as an extended state, the augmented system can be written as:
$\begin{cases}\begin{array}[]{l}\dot{x}_{1i}=x_{2i}\\ \dot{x}_{2i}=F_{i}(x)+G_{i}(x)\dot{q}_{di}+x_{3i}\\ \dot{x}_{3i}=\dot{f}_{i}.\end{array}\end{cases}$ (21)
In many robotic systems, both joint position $q$ and velocity $\dot{q}$ are measurable, enabling the design of second- or third-order ESOs based on (21). With velocity measurements available, a second-order ESO is proposed for the $i$-th joint to estimate $f_{i}$; otherwise, it can be readily extended to a third-order form.
The second-order ESO can be designed as:
$\begin{cases}\begin{array}[]{l}\dot{\hat{x}}_{2i}=F_{i}(x)+G_{i}(x)\dot{q}_{di}+\hat{x}_{3i}+\beta_{1i}(x_{2i}-\hat{x}_{2i})\\ \dot{\hat{x}}_{3i}=\beta_{2i}(x_{2i}-\hat{x}_{2i}),\end{array}\end{cases}$ (22)
where $\hat{x}_{2i}$ and $\hat{x}_{3i}$ are the estimated velocity for $\dot{q}_{i}$ and estimated disturbance for $f_{i}$, respectively. $\beta_{1i}$ and $\beta_{2i}$ are observer gains. The estimation error dynamic is derived from (21) and (22):
$\dot{\vartheta}_{i}=A_{i}\vartheta_{i}+E_{i}\dot{f}_{i},\\$ (23)
where $\vartheta_{i}=\begin{bmatrix}x_{2i}-\hat{x}_{2i}\\ x_{3i}-\hat{x}_{3i}\\ \end{bmatrix}$, $A_{i}=\begin{bmatrix}-\beta_{1i}&1\\ -\beta_{2i}&0\\ \end{bmatrix}$, $E_{i}=\begin{bmatrix}0\\ 1\end{bmatrix}$.
The observer gains are chosen such that the eigenvalues of the matrix $A_{i}$ are placed at $\omega_{o_{i}}$, where $\omega_{o_{i}}$ is the only tunable value known as the observer bandwidth [30].
IV-C Controller Design
$\hat{f}$ can be used to develop a disturbance rejection controller that complements the control law (13):
[…]
where $\dot{q}_{d_{0}}$ is the nominal control from (13), $\dot{q}_{d_{f}}$ rejects disturbance, and they are summed to get $\dot{q}_{d}$.
IV-D Controller Stability Analysis
Assumption 3
There exists a positive known constant $l_{f}$ such that for any $x\in\mathcal{X}$, $u\in\mathcal{U}$, and $t\geq 0$, the following inequality holds:
$\left|\dfrac{\partial f_{i}(t,x,u)}{\partial t}\right|\leq l_{f}.$ (25)
Assumption 3 implies that $f$ is locally Lipschitz continuous with respect to time, and $\dot{f}$ is bounded.
Theorem 1
Given the robotic system (3) with kinematic control input, the ESO (22) with appropriate observer bandwidth $\omega_{o_{i}}$ for each joint, and the nominal control law (24) with appropriate tuned $k_{p}$ and $k_{d}$, we have the following properties: (i): the closed-loop system of (3) is bounded stable if each $\omega_{o_{i}}$ is finite; (ii): the closed-loop system of (3) is asymptotically stable if each $\omega_{o_{i}}$ approaches infinity.
Proof 1
Since $A_{i}$ is Hurwitz, we have the following decomposition [31, Chapter 1.3]:
$e^{A_{i}t}=P_{i}e^{\Lambda_{i}t}P_{i}^{-1},$ (26)
where $P_{i}$ is a matrix whose columns are the eigenvectors of $A_{i}$, and $\Lambda_{i}$} is a diagonal matrix with the corresponding eigenvalues in the diagonal elements.
Therefore, there exists a constant $c_{1_{i}}$ such that [31, Chapter 1.9]
$\left\lVert e^{A_{i}t}\right\rVert\leq\left\lVert P_{i}\right\rVert\left\lVert e^{\Lambda_{i}t}\right\rVert\left\lVert P_{i}^{-1}\right\rVert=c_{1_{i}}\left\lVert e^{\Lambda_{i}t}\right\rVert\leq c_{1_{i}}\left\lVert e^{a_{i}t}\right\rVert,$ (27)
where $a_{i}$ is the $\lambda_{max}(\Lambda_{i})$, the maximum eigenvalue of $\Lambda_{i}$, which, in our case, is $\omega_{o_{i}}$.
The solution to the error dynamics in (23) is given by:
$\vartheta_{i}=e^{A_{i}(t-t_{0})}\vartheta_{i}(t_{0})+\int_{t_{0}}^{t}e^{A_{i}(t-\tau)}E\dot{f}_{i}(\tau)\,d\tau,$ (28)
where $t_{0}$ is the initial time. Substituting (27) into (28), we obtain:
[…]
Since $\dot{f}_{i}$ is bounded, $e^{-\omega_{o_{i}}(t-t_{0})}$ decays exponentially, the estimation error $\left\lVert\vartheta_{i}\right\rVert\ \leq\sigma_{i}$ is bounded. With $\omega_{o_{i}}$ approaching infinity, $\sigma_{i}$ approaches 0.
Thus, the disturbance estimation error satisfies:
$\left\lVert\tilde{f}\right\rVert=\left\lVert f-\hat{f}\right\rVert\leq\sigma.$ (30)
For the $i$-th joint subsystem, substituting the control law (24) into (3) results in the following $i$-th closed-loop subsystem dynamics:
$\ddot{q}_{i}=\ddot{q}_{i}^{\star}+k_{p_{i}}(q_{i}^{\star}-q_{i})+k_{d_{i}}(\dot{q}_{i}^{\star}-\dot{q}_{i})+\tilde{f}_{i},$ (31)
where $\tilde{f}_{i}=f_{i}-\hat{f}_{i}$ represents the disturbance estimation error. The error dynamic can be expressed as:
$\dot{\epsilon}=H\epsilon+\xi,$ (32)
where $\epsilon=\begin{bmatrix}q_{i}-q_{i}^{\star}\\ \dot{q}_{i}-\dot{q}_{i}^{\star}\\ \end{bmatrix}$, $H=\begin{bmatrix}0&1\\ -k_{p_{i}}&-k_{d_{i}}\\ \end{bmatrix}$, and $\xi=\begin{bmatrix}0\\ \tilde{f}_{i}\\ \end{bmatrix}$.
The matrix $H$ can be designed to be Hurwitz. Choose $V=\frac{1}{2}\epsilon^{T}Q\epsilon$ as the Lyapunov function candidate. Its derivative along (32) is given by:
$\dot{V}=\tfrac{1}{2}(\dot{\epsilon}^{T}Q\epsilon+\epsilon^{T}Q\dot{\epsilon})=-\|\epsilon\|^{2}+\xi^{T}Q\epsilon.$ (33)
By applying the Cauchy–Schwarz inequality, we obtain:
[…]
Given that $Q$ is positive definite, $\left\lVert Q\right\rVert\leq\lambda_{max}(Q)$, and the $\xi$ is bounded by $\sigma$, the inequality (34) becomes:
$\begin{array}[]{ll}\dot{V}&\leq-\left\lVert\epsilon\right\rVert^{2}+\sigma\lambda_{max}(Q)\left\lVert\epsilon\right\rVert.\end{array}$ (35)
Next, using the inequality:
$\begin{array}[]{ll}\sigma\lambda_{max}(Q)\left\lVert\epsilon\right\rVert\leq\frac{1}{2}\left\lVert\epsilon\right\rVert^{2}+\frac{1}{2}\sigma^{2}\lambda^{2}_{max}(Q),\end{array}$ (36)
(35) can be rewritten as:
$\begin{array}[]{ll}\dot{V}&\leq-\frac{1}{2}\left\lVert\epsilon\right\rVert^{2}+\frac{1}{2}\sigma^{2}\lambda^{2}_{max}(Q)\\ &\leq-\frac{V}{\lambda_{max}(Q)}+\frac{1}{2}\sigma^{2}\lambda^{2}_{max}(Q).\end{array}$ (37)
Solving the inequality (37) yields:
$\begin{array}[]{ll}{V}&\leq(v(0)-\frac{\kappa}{\eta})e^{-\eta t}+\frac{\kappa}{\eta},\end{array}$ (38)
where $\eta=\frac{1}{\lambda_{max}(Q)}$ and $\kappa=\frac{1}{2}\sigma^{2}\lambda^{2}_{max}(Q)$. 1) It can be concluded from (38) that the states of the closed-loop system are bounded. 2) Moreover, as $\omega_{o}$ approaches infinity, $\sigma$ approaches zero, reducing the disturbance estimation error. Consequently, the inequality (35) is simplified as $\dot{V}\leq-\eta V$, which implies that the system is asymptotically stable.
IV-E Robust CBF for Safe Control
By adding the disturbance term, system (11) becomes:
$\dot{x}=\psi(x)+g(x)u+\bar{g}f,$ (39)
where $f=[f_{1},\cdots,f_{n}]^{T}\in\mathcal{F}\subset\mathbb{R}^{n}$ is a total disturbance vector in each input channel, $\bar{g}=[\mathbf{0}_{n\times n},\mathbf{I}_{n\times n}]^{T}$, $\mathbf{0}_{n\times n}$ and $\mathbf{I}_{n\times n}$ are $n\times n$ zero and identity matrices, respectively. Note that $u$ and $f$ are in the same channel, i.e., the $n\times n$ matrix of the upper part of $g(x)$ is a zero matrix and each row in the lower part of $g(x)$ includes an independent input, so $f$ is a matched disturbance vector.
To define safety sets for system (39), we consider an $r$-times continuously differentiable function $h(x)$, where the relative degrees of $h(x)$ with respect to both $u$ and $f$ are $r$, given that $f$ is matched. A series of functions are defined as follows:
$\begin{array}[]{ll}h_{0}(x)=&h(x),\\ h_{1}(x)=&\dot{h}_{0}(x)+\gamma_{1}h_{0}(x),\\ &\vdots\\ h_{r}(x)=&\dot{h}_{r-1}(x)+\gamma_{r}h_{r-1}(x),\end{array}$ (40)
where $\gamma_{1},\cdots,\gamma_{r}$ are positive constants. The corresponding series of safety sets are
$\begin{array}[]{ll}\mathcal{C}_{0}=&\{x\in\mathbb{R}^{2n}:h_{0}(x)\geq 0\},\\ \mathcal{C}_{1}=&\{x\in\mathbb{R}^{2n}:h_{1}(x)\geq 0\},\\ &\vdots\\ \mathcal{C}_{r}=&\{x\in\mathbb{R}^{2n}:h_{r}(x)\geq 0\}.\end{array}$ (41)
Theorem 2
For system (39), the relative degrees of $h(x)$ with respect to both the input $u$ and the disturbance $f$ are $r$, and $f$ is known. If the initial states satisfy $x(0)\in\mathcal{C}_{s}=\cap_{j=0}^{r}\mathcal{C}_{j}$, then any Lipschitz continuous controller $u(x)\in K_{\text{cbf}}(t,x,f)$ renders the set $\mathcal{C}_{s}$ forward invariant for system (39), where
$\begin{array}[]{r@{}l}K_{\text{cbf}}(t,x,f)\triangleq&\{u\in\mathcal{U}:L_{\psi}^{r}h(x)+L_{g}L_{\psi}^{r-1}h(x)u\\ &+L_{\bar{g}}L_{\psi}^{r-1}h(x)f+\sum\limits_{j=0}^{r-1}k_{j}h^{(j)}(x)\geq 0\},\end{array}$ (42)
where $h^{(j)}(x)$ is the $j$th time derivative of $h(x)$, and $k_{j},j=0,\cdots,r-1$, are the coefficients of polynomial $s^{r}+k_{r-1}s^{r-1}+\cdots+k_{0}$ with roots at $\gamma_{1},\cdots,-\gamma_{r}$.
Proof 2
From (40), $h_{r}$ can be written as
$\begin{array}[]{r@{}l}h_{r}(x)=&(s+\gamma_{r})(s+\gamma_{r-1})\cdots(s+\gamma_{1})h(x)\\ =&s^{r}h(x)+k_{r-1}s^{r-1}h(x)+\cdots+k_{0}h(x),\end{array}$ (43)
where $s\triangleq\frac{d}{dt}$. Since the control input $u$ and disturbance $f$ both have the same relative degree of $r$, $u$ and $f$ do not explicitly show in $h^{(j)}(x),0\leq j\leq r-1$ until
$h^{(r)}(x)=L_{\psi}^{r}h(x)+L_{g}L_{\psi}^{r-1}h(x)u+L_{\bar{g}}L_{\psi}^{r-1}h(x)f.$ (44)
By substituting (44) into (43) and comparing it with (42), the control values in $K_{\text{cbf}}(t,x,f)$ guarantee $h_{r}(x)\geq 0$, $\forall t>0$. From the last equation in (40), we have $\dot{h}_{r-1}(x)+\gamma_{r}h_{r-1}(x)\geq 0$. Then $\dot{h}_{r-1}(x)\geq 0$ for any $x\in\partial\mathcal{C}_{r-1}$. According to Nagumo’s theorem [32], since $x(0)\in\mathcal{C}_{r-1}$, we have $h_{r-1}(x)\geq 0$, $\forall t>0$. Iteratively, $\dot{h}_{0}(x)+\gamma_{1}h_{0}(x)\geq 0$, and we have $h_{0}(x)\geq 0$, $\forall t>0$, as $x(0)\in\mathcal{C}_{0}$. Therefore, the set $\mathcal{C}_{s}$ is forward invariant for system (39). This completes the proof of Theorem 2.
However, the disturbance vector $f$ is not available in practice to enforce the inequality in (42). The estimated disturbance vector $\hat{f}$ obtained in Subsection IV-B is used to devise our robust CBF. Then, (42) can be reformulated as:
$\begin{array}[]{r@{}l}K_{\text{cbf}}(t,x,f,\hat{f})\triangleq&\{u\in\mathcal{U}:L_{\psi}^{r}h(x)+L_{g}L_{\psi}^{r-1}h(x)u\\ &+L_{\bar{g}}L_{\psi}^{r-1}h(x)\hat{f}-L_{\bar{g}}L_{\psi}^{r-1}h(x)(\hat{f}-f)\\ &+\sum\limits_{j=0}^{r-1}k_{j}h^{(j)}(x)\geq 0\}.\end{array}$ (45)
Since the disturbance estimation error $\hat{f}-f$ is unknown in (45), the following assumption is made for the disturbance estimation error bound of ESO.
The estimation error of $\hat{f}_{i}$ is as follows [29]:
$f_{i}(k)-\hat{f}_{i}(k)=p(k)*\Delta f_{i}(k),$ (46)
where * represents convolution and
$\footnotesize\begin{array}[]{l}p(k)=\\ \begin{cases}1&1\leq k\leq r_{i}+1\\ \sum\limits_{\iota=1}^{r_{i}+1}\dfrac{1}{(\iota-1)!}\left(\prod\limits_{j=-\iota+1}^{-1}(k+j)\right)(1-\omega_{o_{i}})^{\iota-1}\omega_{o_{i}}^{k-\iota}&k\geq r_{i}+2.\end{cases}\end{array}$ (47)
For simplicity, let $\prod_{j=0}^{-1}(k+j)=1$. Note that (46) is the disturbance estimation error formulated in the discrete-time domain. $f_{i}(k)=f_{i}(kT_{s})$, $\hat{f}_{i}(k)=\hat{f}_{i}(kT_{s})$, $\Delta f_{i}(k)=f_{i}((k+1)T_{s})-f_{i}(kT_{s})$, $T_{s}$ is the sample time used in calculating the error bound, and $r_{i}$ and $\omega_{o_{i}}$ are the relative degree of $\dot{q}_{i}$ (output measurement) with respect to $f_{i}$ (disturbance as input) in the $i$-th joint subsystem and the poles of the $i$-th ESO in the discrete-time domain, respectively. Readers are referred to our previous work [29] for a complete proof of this error bound.
From the definition of derivative, Assumption 3, and (46), the disturbance estimation error bound of ESO is
$|f_{i}(k)-\hat{f}_{i}(k)|\leq\Gamma_{i}(\omega_{o_{i}},T_{s})=\left(\sum\limits_{k=1}^{\infty}p(k)\right)l_{f}T_{s}.$ (48)
As mentioned in [25], although (48) is derived in the discrete-time domain, the disturbance estimation error bound obtained from it is still the same as that from the continuous-time domain. Since we need to utilize the disturbance estimation error bound in the continuous-time domain, $T_{s}$ should be small, such as 0.1 ms, in our experiment setting. $\omega_{o_{i}}$ in the discrete-time domain needs to be converted from the continuous-time domain through Z-transform.
Theorem 3
Given the system (39) and the ESO in (22) under Assumption 3, any controller $u(x)\in K_{\text{rcbf}}$ renders the set $\mathcal{C}_{s}$ forward invariant for system (39), where
$\begin{array}[]{r@{}l}K_{\text{rcbf}}(t,x,f,\hat{f})\triangleq&\{u\in\mathcal{U}:L_{\psi}^{r}h(x)+L_{g}L_{\psi}^{r-1}h(x)u\\ &+L_{\bar{g}}L_{\psi}^{r-1}h(x)\hat{f}-|L_{\bar{g}}L_{\psi}^{r-1}h(x)|\Gamma(\omega_{o},T_{s})\\ &+\sum\limits_{j=0}^{r-1}k_{j}h^{(j)}(x)\geq 0\},\end{array}$ (49)
$|L_{\bar{g}}L_{\psi}^{r-1}h(x)|$ denotes the absolute value of each element, and $\Gamma(\omega_{o},T_{s})=[\Gamma_{1}(\omega_{o_{1}},T_{s}),\cdots,\Gamma_{n}(\omega_{o_{n}},T_{s})]^{T}$.
Proof 3
From Theorem 2, we need to prove $h_{r}(x)\geq 0$, $\forall t>0$. Substituting (44) and (48) into (43) yields
$\begin{array}[]{r@{}l}h_{r}(x)=&L_{\psi}^{r}h(x)+L_{g}L_{\psi}^{r-1}h(x)u+L_{\bar{g}}L_{\psi}^{r-1}h(x)f\\ &+\sum\limits_{j=0}^{r-1}k_{j}h^{(j)}(x)\\ =&L_{\psi}^{r}h(x)+L_{g}L_{\psi}^{r-1}h(x)u+L_{\bar{g}}L_{\psi}^{r-1}h(x)\hat{f}\\ &-L_{\bar{g}}L_{\psi}^{r-1}h(x)(\hat{f}-f)+\sum\limits_{j=0}^{r-1}k_{j}h^{(j)}(x)\\ \geq&L_{\psi}^{r}h(x)+L_{g}L_{\psi}^{r-1}h(x)u+L_{\bar{g}}L_{\psi}^{r-1}h(x)\hat{f}\\ &-|L_{\bar{g}}L_{\psi}^{r-1}h(x)|\Gamma(\omega_{o},T_{s})+\sum\limits_{j=0}^{r-1}k_{j}h^{(j)}(x)\geq 0.\end{array}$ (50)
Then, we have $h_{r}(x)\geq 0$, $\forall t>0$ as $x(0)\in\mathcal{C}_{r}$. Using the same procedure in Theorem 2, the set $\mathcal{C}_{s}$ is forward invariant for system (39). This completes the proof of Theorem 3.
Remark 2
The key difference between our theoretical results in this work and our preliminary version [25] is that we have extended the robust CBF’s safety guarantee from a relative degree of one to a general case of higher order CBF.
The safety specification requires the end-effector to maintain a non-negative distance from a virtual wall. The safety function is defined as $h(x)=y-y_{0}$, where $y$ and $y_{0}$ denote the positions of the end-effector and the virtual wall, respectively. It can be verified that $h(x)$ has a relative degree of two [26]:
[…]
where a positive constant $\gamma_{1}=\gamma_{2}=\gamma$ is chosen,
$\dot{y}=\left(\frac{\partial\zeta}{\partial q}\right)\dot{q},~\text{and}~\ddot{y}=\frac{d}{dt}\left(\frac{\partial\zeta}{\partial q}\right)\dot{q}+\frac{\partial\zeta}{\partial q}\ddot{q}.$
Combining equations (11), (19), and (51) yields
$\begin{array}[]{ll}h_{2}=&2\gamma\dot{y}+\gamma^{2}(y-y_{0})+\frac{d}{dt}\left(\frac{\partial\zeta}{\partial q}\right)\dot{q}\\ &+\frac{\partial\zeta}{\partial q}\Bigl(-\bar{M}^{-1}(\bar{C}\dot{q}+\bar{G}-\bar{K}_{d}\dot{q}_{d})+f)\Bigr).\end{array}$ (52)
To simultaneously achieve tracking and safe control, the nominal controller for $\dot{q}_{d}$, designed in (24), should be minimally intervened. Thus, the following QP-CBF is constructed:
$\dot{q}_{d_{\text{safe}}}=$ $\operatorname*{arg\,min}_{u\in\mathcal{U}}\|u-\dot{q}_{d}\|^{2}$ (53)
s.t. $h_{i}(x(0))>0,~i=0,1,2$
${L_{\psi}^{2}h(x)}+L_{g}L_{\psi}h(x)u$
$+L_{\bar{g}}L_{\psi}h(x)\hat{f}-{|L_{\bar{g}}L_{\psi}h(x)|}\Gamma(\omega_{o},T_{s})$
$+2\gamma\dot{h}(x)+\gamma^{2}h(x)\geq 0.$
V Hardware Experiments
We use the PUMA 500 robot for hardware experiments, as shown in Fig. 3. It has three main joints and a spherical wrist, providing six degrees of freedom. $q_{1}$, $q_{2}$, and $q_{3}$ correspond to the waist, shoulder, and elbow joints, while $q_{4}$, $q_{5}$, and $q_{6}$ represent the wrist rotation, bend, and flange angles. Since our focus is task-space tracking (i.e., wrist center position), only the first three joints are used. Each joint is driven by a DC motor powered by an Adept Technology MV-19 Power Chassis. The control program is built in MATLAB Simulink, which generates C code and an SDF file, and runs via a dSPACE 1103 for control and data acquisition.
Figure 3: PUMA 500 robot used as a testbed in our research.
The inner-loop controller is configured as $\tau=-K_{p}({q}-{q}_{d})-K_{d}(\dot{q}-\dot{q}_{d})$, where $K_{p}=\mathrm{diag}([18,144,9])$ and $K_{d}=\mathrm{diag}([18,12,9])$. The gains are non-round values due to the actuator driver’s voltage-to-torque conversion ratio. The structure and true parameters remain unknown to our outer-loop controller. The nominal second-order system model (12) is used for the controller. The reference trajectory for joints $q_{1}$, $q_{2}$, and $q_{3}$ is $\begin{bmatrix}0.5\sin(t)&0.25\sin(2t)+0.534&0.25\sin(2t)+\frac{\pi}{2}\end{bmatrix}^{T}$. The tracking performance of the nominal inner-loop controller is illustrated in Fig. 4. Specifically, Fig. 4a depicts the trajectory tracking performance in the joint space, while Fig. 4b shows the trajectory tracking performance in the Cartesian space. The results indicate that the inner-loop controller exhibits poor tracking performance. We will demonstrate the improvement achieved by adding our approach to the tracking and safe control tasks.
We use the nominal model (12) throughout the experiments. A key parameter in the controller design is the control gain in (24), $b_{0}\triangleq\bar{M}^{-1}\bar{K}_{d}$, which depends on the nominal inertia matrix $\bar{M}$ and the inner-loop proportional gain. The nominal inertia $\bar{M}$ is adopted from [33]. Following a common ESO tuning rule, we tune each joint individually by initially setting $b_{0}$ large to ensure stability and then gradually reducing it to improve tracking accuracy. The final nominal control gain is $\mathrm{diag}([20,40,10])$. The observer and control bandwidths are set to $80~\mathrm{rad/s}$ and $10~\mathrm{rad/s}$, respectively, for each joint. We refer readers to [34] for additional instructions and theoretical support on tuning nominal control gain.
(a) Joint space performance
(b) Task space performance
Figure 4: Poor inner-loop PD controller’s performance for $10$ s. (a) Trajectory tracking performance in the joint space. (b) Trajectory tracking performance in the Cartesian space. The inner-loop controller can not track the reference trajectory.
V-A Tracking Control
We compare the tracking performance of ours, with [2]. The reason is that we share the same minimal assumptions about the inner-loop controller, which, to the best of our knowledge, are the most general in the literature.
V-A1 Scenario 1: Normal Case without Payload
The robot tracks a reference trajectory without a payload. The tracking performance of the learning-based adaptive controller [2] and ours is shown in Fig. 5. Fig. 5a compares the reference (dashed) and actual joint trajectories of $q_{1}$, $q_{2}$, and $q_{3}$, where blue and red lines represent our method and [2], respectively. The x-axis denotes time (s), and the y-axis denotes joint angles (rad). Fig. 5b shows the Cartesian trajectories in $x$, $y$, and $z$. The close alignment of the final and reference trajectories indicates that both controllers achieve accurate tracking despite model uncertainties and an imperfect inner loop.
The proposed method achieves tracking performance comparable to the learning-based adaptive controller in the no-payload experiment. As shown in Fig. 6, both approaches yield similarly smooth control signals. Transient responses occur initially, particularly in the second joint (Fig. 5a), due to the offset between the initial and reference positions. The learning-based controller exhibits a longer transient phase (see Figs. 5a and 6). This reflects a typical trade-off in adaptive control between transient duration and tracking accuracy.
(a) Joint space comparison
(b) Task space comparison
Figure 5: Comparative tracking precision analysis: (a) Root Mean squared error (RMSE) in joint space: Proposed ($4.1\times 10^{-3}$, $2.76\times 10^{-2}$, $7.6\times 10^{-3}$) radians vs. [2] ($4.3\times 10^{-3}$, $2.7\times 10^{-2}$, $5.7\times 10^{-3}$) radians for each joint. (b) Cartesian tracking error RMSE: Proposed ($2.3$ mm, $2.4$ mm, $15.5$ mm) vs. [2] ($2.9$ mm, $2.1$ mm, $16.1$ mm) for each joint. Data sampled at 1 kHz over a 10 s trajectory execution.
Figure 6: Inner-loop control command comparison for $10$ s. $u_{1}$, $u_{2}$, and $u_{3}$ are the control inputs (voltages) for joint $q_{1}$, $q_{2}$, and $q_{3}$, respectively.
V-A2 Scenario 2: Robustness Testing with Different Payloads
We evaluate controller robustness with payloads of 1 kg, 1.5 kg, and 2 kg attached to the end-effector. Each payload experiment is repeated five times for ANOVA analysis, with identical controller and observer parameters. Fig. 7a and Fig. 7b present the tracking error ANOVA results for both methods. As shown earlier, our approach achieves a shorter transient time than the learning-based adaptive controller. For clarity, transient data (first 5 s) are excluded, and the plots show steady-state tracking errors (after 5 s).
(a) ANOVA analysis for ours
(b) ANOVA analysis for learning-based adaptive control [2]
Figure 7: Steady-state tracking error comparison under different payloads. (a) ANOVA of ours. (b) ANOVA of the learning-based adaptive control [2]. The proposed method shows consistent tracking across payloads, as indicated by uniform box plots, while the adaptive method exhibits larger error variations, implying reduced robustness. Significance levels: ns ($p>0.05$), * ($p\leq 0.05$), ** ($p\leq 0.01$), *** ($p\leq 0.001$). Each experiment is repeated five times.
As shown in Fig. 7a, the ANOVA results indicate that our method maintains statistically consistent errors across different payloads ($p>0.05$), demonstrating strong robustness, whereas the adaptive method shows significant variations with payload changes ($p<0.05$).
In summary, our approach offers several advantages over the learning-based adaptive control [2]. As shown in Fig. 5, it achieves comparable tracking performance without a payload, lower steady-state error, and stable performance under varying payloads (Fig. 7). It also exhibits a shorter transient response, as seen in the trajectories and control signals during the first second (Figs. 5 and 6). Moreover, our design is simpler than a neural network, requiring fewer tunable parameters and offering greater interpretability. Finally, the proposed disturbance compensation supports both tracking and safe control with state constraints, whereas [2] cannot handle constraints.
V-B Safe Control
It is worth noting that $\gamma$ in (53) critically influences the CBF’s behavior. A larger $\gamma$ drives the trajectory closer to the safety set boundary (more aggressive), while a smaller $\gamma$ keeps it farther away (more conservative). To achieve a safe yet efficient controller, we set $\gamma=10$ in all experiments.
We have conducted experiments in three scenarios: (1) Safety and Robustness Testing with Different Payloads, (2) Safety Testing Subject to External Disturbance, and (3) Safety Testing Subject to Gravity Disturbance. The first scenario will be described in detail, whereas the second and third scenarios are demonstrated in our accompanying video.
A virtual boundary is defined at $y_{0}=\begin{bmatrix}y_{1}&-0.1&y_{3}\end{bmatrix}^{T}$ as a spatial constraint for the end effector. It moves freely along the $x$ and $z$ axes but is constrained along the $y$ axis with a lower limit of 0.1-0.1 m. The system is tested with payloads of $1$, $1.5$, and $2$ kg attached to the end effector.
Three approaches are compared: (1) conventional CBF with a nominal model (see Fig. 8); (2) our robust CBF with a nominal model and an ESO for disturbance estimation, compared to the DOB-CBF [16] (see Fig. 9a); and (3) our robust CBF with a nominal model, ESO for disturbance estimation, and estimation error bound, compared to the DOB-CBF with estimation error bound (see Fig. 9b).
Figure 8: Safety boundary: $\bm{y_{0}}=\begin{bmatrix}\bm{y_{1}}&\bm{-0.1}&\bm{y_{3}}\end{bmatrix}^{T}$. Conventional CBF with nominal model: safety performance with variances under different payloads, with the lowest value of $h$ function at $5.9$ mm.
(a) Estimation error bounds not included
(b) Estimation error bounds included
Figure 9: Safety boundary: $\bm{y_{0}}=\begin{bmatrix}\bm{y_{1}}&\bm{-0.1}&\bm{y_{3}}\end{bmatrix}^{T}$. (a) CBFs with the nominal model and disturbance compensation by ESO (solid lines) and DOB (dashed lines) show consistent (robust) safety performance under different payloads, with the lowest $h$ value being 0.366-0.366 mm (unsafe) for our method and 0.597-0.597 mm (unsafe) for the DOB-CBF. (b) CBFs with the nominal model, disturbance compensation by ESO (solid lines) and DOB (dashed lines), and error bounds show consistent (robust) safety performance under different payloads, with the lowest $h$ value being 26.1 mm (safe and less conservative) for our method and 56.8 mm (safe and more conservative) for the DOB-CBF.
Fig. 8 shows the safety performance of the CBF using only the nominal second-order model (12), which neglects model errors, inner-loop uncertainty, and external disturbances. The safety function $h$ varies under different payloads, revealing limited robustness.
We implement the DOB-CBF [16] for comparison. The DOB-CBF estimates $b_{e}(x,d)=\frac{\partial\zeta}{\partial q}f$ in the CBF, while our method directly estimates $f$ in the original dynamics. Thus, the two observers target different quantities, and their gains are not directly comparable. For fairness, excluding error bounds, we tune both observers’ gains to achieve similar minimum $h$ magnitudes, indicating comparable safety performance. Fig. 9a compares the $h$ functions of the DOB-CBF and the proposed method, showing that:
  1. 1.
    The safety function $h$ exhibits consistent performance under different payloads, indicating robustness of both methods.
  2. 2.
    A slight safety violation occurs due to disturbance estimation error, with minimum values of 0.597-0.597 mm for DOB-CBF and 0.366-0.366 mm for the proposed method, highlighting the need to account for estimation error bounds in both.
We collect experimental data for $\hat{b}_{e}(x,d)$ and $\hat{f}$ from the DOB and ESO, respectively. Their finite differences yield the maximum rates of change, used as estimation error bounds. Fig. 9b shows the resulting $h$ values after incorporating these bounds. Compared with Fig. 9a, we observe that:
  1. 1.
    The trajectories remain safe and consistent across different payloads, demonstrating the robustness of both methods.
  2. 2.
    Incorporating the estimation error bound raises $h$, keeping trajectories strictly within the safety set (positive $h$ values).
  3. 3.
    With error bounds, our method stays closer to the safety boundary than [16], indicating lower conservatism. This may be because [16] estimates $b_{e}(x,d)=\frac{\partial\zeta}{\partial q}f$, which includes a state-dependent Jacobian, making it more sensitive to measurement noise and compounding estimation errors.
VI Conclusion and future work
This paper presents a unified framework that improves tracking performance and ensures safe kinematic control for closed-architecture robotic systems. The proposed method integrates seamlessly into existing cascade control structures as an add-on, enabling robust and safe control under uncertain dynamics, external disturbances, and unknown inner-loop controllers. Stability and safety are theoretically guaranteed, and real-time validation is conducted on the PUMA 500 industrial robot. Future work will extend the framework to visual servoing systems, addressing uncertainties in the vision pipeline.

### v2

@@
 Abstract
-In commercial robotic systems, it is common to encounter a closed inner-loop torque controller that is not user-modifiable.
+In commercial robotic systems, it is common to encounter a closed inner-loop (low-level) torque controller that is not user-modifiable.
 However, the outer-loop controller, which sends kinematic commands such as position or velocity for the inner-loop controller to track, is typically exposed to users.
@@
 This is particularly beneficial when 1) the inner-loop controller is imperfect, unmodifiable, and uncertain; and 2) the dynamic model exhibits significant uncertainty.
-Stability analysis, formal safety guarantee proof, and hardware experiments with a PUMA robotic manipulator are presented.
+Stability analysis, formal safety guarantee proof, simulations, and hardware experiments with a PUMA robotic manipulator are presented.
 Our solution demonstrates superior performance in terms of simplicity of implementation, robustness, tracking precision, and safety compared to the state of the art.
@@
 I Introduction
-Robotic systems typically use a hierarchical architecture combining perception, planning, and low-level control, which improves modularity and reliability.
-In the control layer, a kinematic controller (outer loop) and a dynamic controller (inner loop) are commonly used.
-Since tuning the inner loop requires expertise, it is usually hidden from users for safety, forming a closed architecture [1, 2, 3, 4].
-Users can thus only design the kinematic controller, sending position or velocity commands.
-However, wear, modifications, or payload changes can alter dynamics and degrade inner-loop performance [5].
-This work addresses the challenge of achieving unified high-performance and safe tracking control using only the kinematic controller, despite an imperfect, unmodifiable inner loop, and uncertain dynamics.
+Robotic systems often employ hierarchical software design, stacking perception, decision-making, planning, and low-level control.
+Such modularity is particularly beneficial for troubleshooting and improving the reliability of robotic systems.
+For example, in the control block, a combination of a kinematic controller (outer-loop controller) and a dynamic controller (inner-loop controller) is commonly seen in various robots.
+However, because tuning the inner-loop controller requires expert knowledge, this component is typically not exposed to users due to product safety considerations, a practice referred to as closed architecture in the literature [wang2019dynamic, ahanda2022adaptive, ahanda2022task, khan2024control].
+In other words, users are only allowed to design the kinematic controller, sending position or velocity for the inner-loop controller to track.
+Additionally, mechanical parts can wear out over time, and users may modify the robot by adding or replacing parts or attaching payloads, which can lead to changes in dynamics and potentially render the inner-loop controller suboptimal [liu2022advancements].
+The fundamental challenge we address in this article is how to achieve high-performance tracking control and safe control in a unified framework for the entire dynamic system using only the kinematic controller, while the inner-loop controller is imperfect and unmodifiable, and both the dynamic model and the inner-loop controller exhibit significant uncertainty.
 Fig.
@@
 Figure 1: Trajectory tracking of the PUMA 500 for an infinity-shaped reference.
+The trajectory using a poor inner-loop controller is shown in red, while the improved tracking performance (ours) is shown in blue.
+To bear in mind the two control objectives: high-performance tracking and safe control, we briefly survey the literature in the categories of adaptive control and reduced-order model-based barrier functions to illustrate our current position.
 A detailed literature review can be found in the subsequent section.
-Existing works typically rely on the following assumptions: 1) The inner-loop controller is known and accessible to the user; 2) The inner-loop controller is known but not accessible to the user; 3) The inner-loop controller has perfect tracking for outer-loop commands.
+[wang2019dynamic] presents a class of adaptive outer-loop control schemes that generate joint velocity and position commands for the inner-loop controller with certain proportional-integral (PI) or proportional-integral-derivative (PID) structures.
+[ahanda2022adaptive, khan2024control] relaxes the inner-loop controller’s structure and employs neural networks-based adaptive control to approximate the inner-loop controller’s dynamics and both kinematic and dynamic parameters.
+These works only consider tracking control without safety functions, such as state constraints.
+Additional limitations related to assumptions, implementation complexity, and performance will be discussed in the literature review and experimental comparisons.
+Control barrier functions (CBFs) have emerged as a powerful tool for designing controllers that guarantee safety [ames2016control, ames2019control].
+However, most existing CBF approaches rely on high-fidelity dynamic models.
+Reduced-order model-based CBFs [molnar2021model] demonstrate robust obstacle avoidance using only safe velocities, without requiring high-fidelity dynamic models.
+However, [molnar2021model] assumes perfect inner-loop tracking performance.
+[singletary2021safety] and [molnar2023safety] propose CBFs with kinetic energy and input-to-state safe backup set methods, respectively, reducing model dependence and allowing for some model uncertainty.
+However, they still rely on prior knowledge of numerous dynamic model parameters and a known-structure inner-loop controller.
+To sum up, existing works typically rely on the following assumptions: 1) The inner-loop controller is known and accessible to the user; 2) The inner-loop controller is known but not accessible to the user; 3) The inner-loop controller has perfect tracking for outer-loop commands.
 Our goal is to minimize reliance on knowledge of the dynamic model and the inner-loop controller, while still guaranteeing tracking and safe control using only an outer-loop kinematic controller.
+We develop a novel add-on at the kinematic control layer, by combining extended state observer (ESO)-based disturbance rejection control and robust CBF, called ESOR-QP [chen2023robust] for high-performance tracking and provably safe control.
+Distinct from the aforementioned methods, our method treats the imperfect inner-loop controller, unmodeled dynamics, and external disturbance as a total disturbance, estimated by ESO, and then sends it to both the disturbance rejection controller and the robust CBF.
+Note that our solution is not simply a merge of two functionalities in tracking and safe control.
+Instead, a well-estimated total disturbance is naturally used as compensation for disturbance rejection and safe control.
+The so-called compensation has dual benefits in these two components.
+For trajectory tracking in task space, the tracking controller rejects the total disturbance for high-performance trajectory tracking.
+For safe control to avoid collisions, the estimated disturbance is integrated into a robust constraint in the CBF-based optimal control.
 As shown in Fig.
-2, existing robotic systems typically send desired position ($q^{*}$) or velocity ($\dot{q}^{*}$) commands to the inner-loop controller.
-Our solution, inserted between them (gray box), comprises a disturbance-rejection tracking controller, an extended state observer (ESO), and an extended state observer-based robust quadratic programming (ESOR-QP) module.
-The ESO estimates the total disturbance $\hat{f}$ using the joint angle $q$ and the safe kinematic control $q_{d,\text{safe}}$ or $\dot{q}_{d,\text{safe}}$.
-This estimate is compensated in the tracking controller to produce nominal commands $q_{d}$ or $\dot{q}_{d}$, which are monitored by the ESOR-QP to enforce robust safety constraints.
+2, our developed outer-loop controller is composed of a disturbance rejection tracking controller, an ESOR-QP module, and an ESO module.
+First, the ESO uses output measurement $q$ (joint angle) and final safe kinematic control $q_{d_{safe}}$ (position control) or $\dot{q}_{d_{safe}}$ (velocity control) to estimate the total disturbance $\hat{f}$.
+Second, the estimated total disturbance $\hat{f}$ is rejected in the tracking controller to generate outer-loop kinematic control commands $q_{d}$ or $\dot{q}_{d}$.
+Third, our robust CBF, ESOR-QP, continuously monitors the nominal control $q_{d}$ or $\dot{q}_{d}$ and intervenes if the nominal control is deemed unsafe.
+Such a safety filter relies on robust safety constraints integrating the nominal model as well as the estimated disturbance $\hat{f}$.
 The contributions of our work are summarized as follows:
-  • We innovatively formulate the kinematic control of closed-architecture robotic manipulator systems as a disturbance rejection problem, requiring only a practical assumption on the structure, Jacobian, and nominal inertia matrix.
-  • We develop a robust control barrier function (CBF) framework based on an extended state observer, enabling safe control of general nonlinear affine systems with high relative degrees under disturbances.
-To the best of our knowledge, this is the first work achieving safe kinematic control of closed-architecture robots under largely unknown uncertainties.
-  • As shown in Fig.
-2, our method is implemented as a standalone add-on module that enhances existing robotic manipulators without modifying the inner-loop controller.
-Real-time control at 1 kHz is demonstrated on a PUMA 500 manipulator.
+  • System-level contribution: We propose a unified control framework that enhances tracking performance and safety using only a kinematic controller.
+Our approach does not require access to the inner-loop controller and remains effective in the presence of model parameter uncertainties, a largely unknown inner-loop controller, and external disturbances from the environment.
+  • Theoretical contribution: We develop a robust CBF based on an extended state observer, enabling safe control of general nonlinear affine systems with high-order relative degrees under disturbances.
+  • Practical contribution: As illustrated in Fig.
+2, our solution is implemented as a standalone add-on component that can be integrated into existing legacy robotic manipulators to improve performance without modifying the inner-loop controller.
+Real-time control at 1 kHz has been implemented on a PUMA 500 robotic manipulator.
 Figure 2: The proposed framework in this paper, is marked as the Outer-loop Controller.
 The total estimated disturbance $\hat{f}$ is used for both disturbance compensation and safe control using ESOR-QP.
+We have made the following substantial theoretical and experimental improvements over our preliminary conference version [chen2023robust]:
+  1. 1.
+    Theory: We have extended our robust CBF for a system with relative degree one to accommodate arbitrary relative degrees, i.e., higher-order robust CBFs.
+Our work provides rigorous proof, complementing [chen2023robust], for the forward invariance property of the safety set.
+  2. 2.
+    Application: We demonstrate the efficacy of our approach through validation in real-world hardware experiments on a nontrivial robotic manipulator, addressing the particular challenge of a closed architecture.
+In contrast, our previous work was validated on simpler robotic systems in simulation, including a cruise control system with 1D control and a self-balancing scooter with 2D control.
 The rest of the paper is organized as follows: Section II reviews the related work.
@@
 Section IV presents the proposed framework.
-Hardware experiments are reported in Section V.
-Concluding remarks are in Section VI.
+Simulation and hardware experiments are reported in Section V and Section VI, respectively.
+Concluding remarks are in Section VII.
 II Related Work
+In this section, we will briefly review the most important related work on kinematic control of closed architecture and safe control.
 An overview of the comparison between our work and the most related works is presented in Table I.
-II-A Adaptive Control for Closed Architecture
-Learning-based Adaptive Control: [2, 3, 4] employ neural network–based outer-loop adaptation to approximate the kinematics, dynamics, and inner-loop controller.
-As summarized in Table I, they make the weakest assumption about the inner loop, identical to ours, where $\Psi(\cdot)$ denotes an unknown structure possibly containing proportional, integral, or state-dependent terms.
-However, they rely on numerous tuning parameters and limited interpretability.
-Conventional Adaptive Control: [1] introduces dynamic modularity with an adaptive controller under a moderate assumption of a known PI/PID inner-loop structure.
-These studies address tracking control only, without safety considerations.
-In our experiments, we compare tracking performance with the learning-based adaptive control due to the shared inner-loop assumption.
+A vast body of research on adaptive [slotine1987adaptive] and robust [slotine1985robust] torque control has been developed over the past several decades to address various uncertainties in robotic manipulators.
+However, we do not extensively cover this line of work in our literature review due to the unique nature of our problem, which focuses on a robotic manipulator with a closed-architecture system.
+A direct comparison between torque control and kinematic control is not fair, as torque control—implemented in the inner-loop layer of a cascaded control structure—is inherently more responsive than outer-loop approaches.
+Instead, we aim for our detailed formulation and proposed solution to motivate the adaptation and redesign of torque control methodologies to better accommodate our specific context.
+Furthermore, the challenge of dealing with unknown inner-loop controllers remains a promising direction for future research across all existing approaches, including both ours and adaptive control methods.
+II-A Adaptive Control
+Learning-based Adaptive Control: [ahanda2022adaptive, ahanda2022task, khan2024control] use adaptive outer-loop controller based on approximated kinematics, dynamics, and inner-loop controller using a neural network.
+As shown in Table I, these methods make the weakest assumption regarding the inner-loop controller, the same as our work: the $\Psi(\cdot)$ term represents an unknown structure that may include proportional, integral, and state-dependent compensation components.
+However, these approaches involve unexplainable learning processes (e.g., a large number of tuning parameters and troubleshooting difficulties).
+Conventional Adaptive Control:[wang2019dynamic] proposes dynamic modularity, which employs an adaptive control method with a moderate assumption (see Table I): a known structure of the inner-loop controller, either PI or PID.
+Note that these works focus only on tracking control tasks without considering safety.
+In our experiment, we will show a comparison of tracking performance with the learning-based adaptive control, since we share the same assumption of the inner-loop controller.
 TABLE I: Overview of our work’s position in the literature
@@
 $\tau=-K_{d}(\dot{q}-\dot{q}_{d})+\Psi(q,\dot{q},q_{d},\int_{0}^{t}q\,dt,\int_{0}^{t}q_{d}\,dt)$
-[2, 3, 4]
+[ahanda2022adaptive, ahanda2022task, khan2024control]
 Conventional
@@
 Moderate, known PID/PI structure
-[1]
+[wang2019dynamic]
 Reduced-order CBF -
@@
 bounded error
-[6, 7, 8]
-Robust CBF - - [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]
+[molnar2021model, singletary2021safety, molnar2023safety]
+Robust CBF - - [mrdjan2018robust, kolathaya2019input, alan2021safe, buch2022robust, nguyen2022robust, alan2023parameterized, wang2023disturbance, dacs2022robust, alan2023disturbance, zhao2020adaptive, sun2024safety, das2025robust, wang2024safety, cao2024safety, zhang2024eso, wang2023composite]
 Ours
@@
 II-B Reduced-order model-based CBF
-Many CBF-based safety approaches rely on high-fidelity dynamic models with torque control, yet obtaining such models remains challenging.
-Reduced-order or kinematic CBFs have been introduced to ensure safety with lower-fidelity models.
-For instance, [6] guarantees safety via safe velocity control but assumes a perfect inner-loop controller capable of instant tracking.
-Similarly, [7, 8] employ kinetic energy–based CBFs that offer partial robustness to model uncertainty but still depend on numerous dynamic parameters and a known PD-type inner loop.
-In contrast, our work addresses closed-architecture robots with minimal knowledge of the inner-loop controller and significant model uncertainty.
+To ensure the safety of robotic systems, many efforts using CBF theory focus on high-fidelity dynamic models with torque control.
+However, obtaining an accurate high-fidelity model is a fundamental challenge.
+Reduced-order model-based CBFs have been proposed to provide safety guarantees with a kinematic controller, even when using a low-fidelity model.
+[molnar2021model] aims to ensure safety using safe velocity without relying on a high-fidelity dynamical model.
+However, their safety guarantee is based on a strong assumption of a perfect inner-loop controller: any outer-loop velocity command can be tracked quickly.
+[singletary2021safety, molnar2023safety] propose kinetic energy-based CBFs to ensure safety at the kinematic control level, allowing for some robustness to model uncertainty, but still requiring many dynamic model parameters and a known inner-loop controller with a specific PD structure.
+The fundamental difference between this line of research and our work is that we focus on a closed architecture with minimal available information about the inner-loop controller while allowing the dynamic model to exhibit significant uncertainty.
+As shown in Table I, due to the different assumptions, it is difficult to compare the safety performance of these works with ours.
 II-C Robust CBF
-Robust CBFs have recently been developed to handle model uncertainties and external disturbances [9, 10, 11, 12, 13, 14, 15, 17, 19, 20, 21, 25, 18, 16, 23, 24, 22].
-Among them, disturbance observer–based robust CBFs (DOB-CBFs) [25, 18, 16, 23, 24, 22, 15, 17, 19, 20, 21] enable active disturbance estimation without requiring worst-case bounds.
-Unlike these methods where estimation error bounds are absent or loosely derived via Lyapunov analysis, our approach obtains them directly from the observer error dynamics, resulting in tighter and less conservative bounds.
-Experimental comparisons with state-of-the-art methods (e.g., [16]) further demonstrate this advantage.
-In contrast to our conference version [25] that only considered relative-degree-one systems, this work provides a generalized framework with rigorous proofs guaranteeing the forward invariance of the safety set.
+The original CBF [ames2016control] relies on a precise dynamic model to ensure that the system’s trajectory remains within a forward-invariant safety set, guaranteeing safety.
+To address this fundamental limitation in the presence of model uncertainty, robust CBFs have been proposed in recent years.
+These methods recover formal safety guarantees by incorporating additional terms in the constraints of CBF formulation to compensate for model uncertainties and external disturbances [mrdjan2018robust, kolathaya2019input, alan2021safe, buch2022robust, nguyen2022robust, alan2023parameterized, wang2023disturbance, alan2023disturbance, sun2024safety, das2025robust, wang2024safety, chen2023robust, zhao2020adaptive, dacs2022robust, zhang2024eso, wang2023composite, cao2024safety].
+Among these methods, disturbance observer-based robust CBF (DOB-CBF) [chen2023robust, zhao2020adaptive, dacs2022robust, zhang2024eso, wang2023composite, cao2024safety, wang2023disturbance, alan2023disturbance, sun2024safety, das2025robust, wang2024safety] is one of the most popular approaches for handling uncertainties in dynamic models, as it allows for active disturbance estimation without requiring the worst-case disturbance bound [sariyildiz2020disturbance, sadeghian2013task].
+As a result, DOB-CBF approaches, such as [wang2023disturbance], often exhibit less conservatism compared to the robust CBF methods that rely on worst-case disturbance assumptions [nguyen2022robust].
+In [dacs2022robust], a DOB is designed to estimate the disturbance within the CBF dynamics, rather than the original system’s disturbance, to guarantee safety.
+To further enhance robustness and reduce conservativeness, [alan2023disturbance] integrates the DOB from [dacs2022robust] with the concept of tunable input-to-state safe control.
+Instead of using a disturbance observer to estimate disturbances within the CBF dynamics, [zhao2020adaptive] introduces an adaptive law to estimate disturbances in the original system dynamics.
+A sampled-data safe control strategy combined with a DOB is proposed in [sun2024safety] to address uncertainties caused by both intersample time intervals and disturbances.
+In [das2025robust], the original high-order CBF-QP is reformulated as a second-order cone program for systems with a specific structure, where the relative degree differences between the control inputs and disturbances are all less than or equal to 1.
+To handle unmatched disturbances in general nonlinear systems with disturbances, [wang2024safety] employs a DOB to estimate disturbances and the bounds of their higher-order derivatives, which are challenging to obtain in practice.
+In [cao2024safety], a novel safety-critical generalized predictive control method is introduced for speed regulation of a permanent magnet synchronous motor with current and voltage constraints, utilizing fixed-time sliding-mode disturbance observers to estimate disturbances.
+These approaches assume access to all true states.
+However, in practice, only measurements are often available.
+In such cases, state uncertainty needs to be accounted for, e.g., by incorporating an ESO within the CBF framework, as seen in [chen2023robust, zhang2024eso, wang2023composite].
+Note that in these previous works, the estimation error bounds of disturbances are often relaxed using Lyapunov methods, whereas in our approach, they are obtained directly from the solution of the observer’s error dynamics.
+As a result, our method tends to be less conservative.
+We will provide more experimental comparisons on the tightness of the error bounds with state-of-the-art methods, such as [dacs2022robust].
 III Preliminary and Problem Formulation
+In this section, we will briefly go through the necessary preliminary, then introduce our problem formulation.
 III-A System Dynamics
@@
 The output $y\in\mathbb{R}^{m}$, and $\zeta:\mathbb{R}^{n}\to\ \mathbb{R}^{m}$ is the mapping from joint space to task space.
-The most common inner-loop controller for robot manipulators typically consists of a PID controller or PD controller, often combined with dynamic compensation, such as gravity.
-For stability, most inner-loop controllers include a derivative term for damping, $(\dot{q}-\dot{q}_{d})$, where $\dot{q}_{d}$ is a joint velocity command.
+The most common inner-loop controller for robot manipulators typically consists of a PID controller or PD controller, often combined with dynamic compensation, such as gravity [kelly2005control].
+For stability, most inner-loop controllers include a derivative term, $(\dot{q}-\dot{q}_{d})$, where $\dot{q}_{d}$ is a joint velocity command.
 Assumption 1
@@
 $\tau=-K_{d}(\dot{q}-\dot{q}_{d})+\Psi(q,\dot{q},q_{d},\int_{0}^{t}q\,dt,\int_{0}^{t}q_{d}\,dt),$ (2)
-with $K_{d}$ a proportional control gain, and $\Psi(\cdot)$ is a general, unknown function that may include proportional, integral, derivative, and dynamic compensation terms.
-To the best of our knowledge, this assumption is the most general in the literature (see Table I), requiring minimal knowledge of the inner-loop controller, similar to [2], and more general than [1].
+with $K_{d}$ a proportional control gain, and $\Psi(\cdot)$ is a general, unknown function that may include proportional, integral, and dynamic compensation terms.
+To the best of our knowledge, this assumption is the most general in the literature (see Table I), requiring minimal knowledge of the inner-loop controller, similar to [ahanda2022adaptive], and more general than [wang2019dynamic].
 Assumption 2
 The kinematics of the system are assumed to be known for control in task space.
-The system dynamic model combining (1) and (2) is:
-$\begin{split}\ddot{q}&={M}^{-1}\Bigl(-{C}\dot{q}-{G}-F_{r}+{J}^{T}F_{ext}-{K}_{d}\dot{q}+\Psi(\cdot)+{K}_{d}\dot{q}_{d}\Bigr)\end{split}.$ (3)
-We have removed torque control, as it is not accessible in a closed architecture.
-$\dot{q}_{d}$ is the kinematic control to be manipulated in the outer-loop layer, and its control law will be designed.
-Another reason for choosing $\dot{q}_{d}$ instead of $q_{d}$ as the control input is to avoid differentiation-induced noise.
+This assumption is mild in practice, as the geometric parameters of a robotic arm are easy to measure.
+Then, the system dynamic model combining (1) and (2) can be expressed as:
+$\begin{split}\ddot{q}&={M}^{-1}\Bigl(-{C}\dot{q}-{G}-F_{r}+{J}^{T}F_{ext}\\ &-{K}_{d}\dot{q}+\Psi(\cdot)+{K}_{d}\dot{q}_{d}\Bigr)\end{split}.$ (3)
+Note that we have removed torque control, as it is not accessible in a closed architecture.
+$\dot{q}_{d}$ is the kinematic control that will be manipulated in the outer-loop layer, and its control law will be designed.
+Another reason for selecting $\dot{q}_{d}$ as the control input, rather than $q_{d}$, is to avoid the need for differentiation, which could introduce significant noise in practical applications.
 III-B Robot Kinematic Control with Known Model
-Neglecting $F_{r}$ and $F_{ext}$ in (1), the exact system dynamics is used to design the control law:
+If the friction term, $F_{r}$, and the external disturbance, $F_{ext}$, are neglected in (1), the exact system dynamics can be leveraged to design a control law:
 $\begin{split}\dot{q}_{d}&={K}_{d}^{-1}\Bigl({M}u_{0}+{C}\dot{q}+{G}+{K}_{d}\dot{q}-\Psi(\cdot)\Bigr)\end{split},$ (4)
@@
 $k_{p}\in\mathbb{R}^{n\times n}$ and $k_{d}\in\mathbb{R}^{n\times n}$ are the proportional and derivative gain matrices, respectively.
-$\ddot{q}^{\star}$ serves only as a feedforward term, while $u_{0}$ is a position controller.
+Note that $\ddot{q}^{\star}$ serves only as a feedforward term, while $u_{0}$ remains a position controller.
 By substituting (4) and (5) into (3), we get:
 $\ddot{q}=\ddot{q}^{\star}+k_{p}(q^{\star}-q)+k_{d}(\dot{q}^{\star}-\dot{q}).$ (6)
-Proper $k_{p}$ and $k_{d}$ make q{q} converge to ${q}^{\star}$ asymptotically.
+By properly selecting $k_{p}$ and $k_{d}$, q{q} will converge to ${q}^{\star}$ asymptotically.
 III-C Control Barrier Functions
-The system (3) can be rewritten as a nonlinear control-affine system when ignoring the disturbance vector $F_{ext}$:
+The robotic system (3) can be rewritten as a nonlinear control-affine system when ignoring the disturbance vector $F_{ext}$:
 $\dot{x}=\psi(x)+g(x)u,$ (7)
 where $x\triangleq[q,\dot{q}]^{T}\in\mathcal{X}\subset\mathbb{R}^{2n}$, $\psi:\mathbb{R}^{2n}\rightarrow\mathbb{R}^{2n}$ and $g:\mathbb{R}^{2n}\rightarrow\mathbb{R}^{2n\times n}$ are Lipschitz continuous, and $u\triangleq\dot{q}_{d}\in\mathcal{U}\subset\mathbb{R}^{n}$ is a control input vector.
-The safety of system (7) can be guaranteed using a safety set.
+The safety of system (7) can be guaranteed using the concept of a safety set.
 A set $\mathcal{C}$ is considered to be a safety set if it is forward invariant in the state space $\mathcal{X}$, i.e., for system (7) if solutions for some $u\in\mathcal{U}$ starting at any initial safe condition $x(0)\in\mathcal{C}$ satisfy $x(t)\in\mathcal{C}$, $\forall t\geq 0$.
@@
 Definition 1
-(Exponential CBF (ECBF) [26, 27]) Consider system (7) with relative degree $r$ for an $r$-times continuously differentiable function $h$, i.e., $L_{g}L_{\psi}h(x)=\cdots=L_{g}L_{\psi}^{r-2}h(x)=0$ and $L_{g}L_{\psi}^{r-1}h(x)\neq 0,\forall x\in\mathcal{C}$.
+(CBF [ames2016control]) A continuous and differentiable function $h:\mathbb{R}^{2n}\rightarrow\mathbb{R}$ is called a CBF for the system (7) with relative degree 1, if
+$\sup\limits_{u\in\mathcal{U}}\left(L_{\psi}h(x)+L_{g}h(x)u\right)\geq-\beta(h(x))$ (9)
+for all $t\geq 0$ and $x\in\mathcal{X}$, where $L_{\psi}h(x)\triangleq\frac{\partial h(x)}{\partial x}\psi(x)$, $L_{g}h(x)\triangleq\frac{\partial h(x)}{\partial x}g(x)$, and $\beta(s)$ is an extended class $\mathcal{K}$ function [^1] A continuous function: $\beta:[-b,a)\rightarrow[-\infty,\infty)$ for some $a>0,b>0$ is said to belong to extended class $\mathcal{K}$, if it is strictly increasing and $\beta(0)=0$..
+The set of control signals ensuring the system safety is defined as:
+$\begin{array}[]{r@{}l}K_{\text{cbf}}(t,x)\triangleq\{&u\in\mathcal{U}:L_{\psi}h(x)\\ &+L_{g}h(x)u\geq-\beta(h(x))\}.\end{array}$ (10)
+For a system with an arbitrary relative degree, we need the following higher-order CBF.
+Definition 2
+(Exponential CBF (ECBF) [nguyen2016exponential, xiao2019control]) Consider system (7) with relative degree $r$ for an $r$-times continuously differentiable function $h$, i.e., $L_{g}L_{\psi}h(x)=\cdots=L_{g}L_{\psi}^{r-2}h(x)=0$ and $L_{g}L_{\psi}^{r-1}h(x)\neq 0,\forall x\in\mathcal{C}$.
 $h(x)$ is an ECBF if there exists a row vector $K_{a}\in\mathbb{R}^{r}$ satisfying $\forall x\in\mathcal{C}$
 $\sup\limits_{u\in\mathcal{U}}\big(L_{\psi}^{r}h(x)+L_{g}L_{\psi}^{r-1}h(x)u\big)\geq-K_{a}\eta_{b}(x),$ (11)
-where $\eta_{b}(x)=[h(x),\dot{h}(x),\cdots,h^{(r-1)}(x)]^{T}$, $K_{a}=[k_{1},\cdots,k_{r}]$, and the values of $k_{1},\cdots,k_{r}$ satisfy specific properties given in [26, 27].
-To guarantee safety, the control problem is formulated as a quadratic program (QP) with a CBF as a hard constraint [28]:
+where $\eta_{b}(x)=[h(x),\dot{h}(x),\cdots,{h^{(r-1)}(x)}]^{T}$, $K_{a}=[k_{1},\cdots,k_{r}]$, and the values of $k_{1},\cdots,k_{r}$ satisfy specific properties given in [nguyen2016exponential, xiao2019control].
+In order to leverage a CBF to guarantee safety, the control problem is formulated as a quadratic program (QP) with a CBF as a hard constraint. The QP formulation is as follows [ames2016control]:
 $\begin{split}u^{*}(x)&=\operatorname*{arg\,min}_{u\in U}\|u-k(x)\|^{2}\\ \text{s.t.}\quad L_{\psi}^{r}h(x)&+L_{g}L_{\psi}^{r-1}h(x)u\geq-K_{a}\eta_{b}(x),\end{split}$ (12)
 where $k(x)$ is a nominal control law.
-Research Objective: Given the system in (3), the nominal dynamics with limited knowledge of the system and inner-loop controller are defined as
-$\ddot{q}=\bar{M}^{-1}\!\left(-\bar{C}\dot{q}-\bar{G}+\bar{K}_{d}\dot{q}_{d}\right),$ (11)
-where $\bar{M}$, $\bar{C}$, and $\bar{G}$ are nominal parameters, and $\bar{K}_{d}$ is the nominal inner-loop gain.
-The goal is to design a $\dot{q}_{d}$ controller that achieves accurate trajectory tracking and task-space safety (e.g., collision avoidance) under an unknown inner-loop controller, uncertain dynamics, and external disturbances.
+Research Objective: Given the system described in (3), we have the following nominal dynamic system with minimal knowledge of the system and the inner-loop controller:
+$\ddot{q}=\bar{M}^{-1}\Bigl(-\bar{C}\dot{q}-\bar{G}+\bar{K}_{d}\dot{q}_{d}\Bigr),$ (13)
+where $\bar{M}$, $\bar{C}$, and $\bar{G}$ are the nominal model parameters, with $\bar{K}_{d}$ representing the nominal gain of the inner-loop controller.
+The objective is to develop a controller for $\dot{q}_{d}$ that enables the system to accurately follow the reference trajectory and guarantee safety, such as collision avoidance in task space.
+This control design must address significant challenges, including: (i) an imperfect and largely unknown inner-loop controller, (ii) uncertain system dynamics, and (iii) the presence of external disturbances.
 Remark 1
@@
 $\ddot{q}=\bar{M}^{-1}\bar{K}_{d}\dot{q}_{d}.$ (14)
-We use the nominal model (11) for general mathematical derivation, whereas the reduced nominal model (12) is employed in hardware experiments to evaluate the tracking performance and safety under limited information.
+We use the nominal model (14) for all the simulations and experiments in this paper.
 IV Proposed Framework
+This section details the controller design, along with formal proofs of stability and safety guarantee.
 IV-A Robot Kinematic Control with Nominal Model
-The control law in (4) is ideal, assuming known dynamics and an inner-loop controller. To handle uncertainties, we propose the following practical law based on nominal dynamics:
+The control law in (4) is ideal due to its reliance on the dynamic model and the inner-loop controller.
+To address uncertainties, we propose the following practical control law based on nominal values:
 $\begin{split}\dot{q}_{d}&=\bar{K}_{d}^{-1}\Bigl(\bar{M}u_{0}+\bar{C}\dot{q}+\bar{G}\Bigr).\end{split}$ (15)
-The following part shows the dynamics of the close-loop system with our nominal control law (13). We define the parametric discrepancies as:
+The following part will show the resulting system with our nominal control law (15). First, we define the parametric discrepancies as:
 $\begin{split}\Delta K_{d}&\triangleq K_{d}\bar{K}_{d}^{-1},\\ \Delta{M}&\triangleq\Delta K_{d}\bar{M}-{M},\\ \Delta{C}&\triangleq\Delta K_{d}\bar{C}-{C},\\ \Delta{G}&\triangleq\Delta K_{d}\bar{G}-{G}.\end{split}$ (16)
@@
 $\begin{split}&\Bigl(\Delta K_{d}\bar{M}-\Delta{M}\Bigr)\ddot{q}+\Bigl(\Delta K_{d}\bar{C}-\Delta{C}\Bigr)\dot{q}\\ &+\Bigl(\Delta K_{d}\bar{G}-\Delta{G}\Bigr)+F_{r}.\end{split}$ (19)
-The resulting close-loop system is:
+The resulting system is:
 $\begin{split}\ddot{q}=\ddot{q}^{\star}+k_{p}(q^{\star}-q)+k_{d}(\dot{q}^{\star}-\dot{q})+f.\end{split}$ (20)
@@
 $\begin{split}f=(\Delta K_{d}\bar{M})^{-1}\Bigl(&\Delta{M}\ddot{q}+\Delta{C}{\dot{q}}+\Delta{G}-F_{r}\\ &+\Psi(\cdot)+{J}^{T}F_{ext}-K_{d}\dot{q}\Bigr).\end{split}$ (21)
-Let $f$ denote the total disturbance, including internal uncertainties and external disturbances.
-Compared with (6), $f$ degrades system performance.
-To counter this, an ESO [29] is employed to estimate $f$ for disturbance rejection.
+We denote $f$ as the total disturbance, composed of internal dynamic uncertainty and external disturbances.
+Compared with (6), the presence of $f$ degrades the system performance.
+To mitigate this, we employ the ESO [chen2022relationship] to estimate $f$ for further disturbance rejection control.
 IV-B Extended State Observer Design
@@
 $\begin{cases}\begin{array}[]{l}\dot{x}_{1i}=x_{2i}\\ \dot{x}_{2i}=F_{i}(x)+G_{i}(x)\dot{q}_{di}+x_{3i}\\ \dot{x}_{3i}=\dot{f}_{i}.\end{array}\end{cases}$ (23)
-In many robotic systems, both joint position $q$ and velocity $\dot{q}$ are measurable, enabling the design of second- or third-order ESOs based on (21).
-With velocity measurements available, a second-order ESO is proposed for the $i$-th joint to estimate $f_{i}$; otherwise, it can be readily extended to a third-order form.
+In many robotic systems, both the joint position $q$ and the velocity $\dot{q}$ are measurable.
+This allows the design of second-order and third-order ESOs based on the augmented system (23).
+Given the availability of velocity measurements, we propose a second-order ESO for the $i$-th joint subsystem to estimate the disturbance $f_{i}$.
+If velocity measurements are unavailable, the observer can be easily modified to accommodate a third-order form.
 The second-order ESO can be designed as:
@@
 where $\hat{x}_{2i}$ and $\hat{x}_{3i}$ are the estimated velocity for $\dot{q}_{i}$ and estimated disturbance for $f_{i}$, respectively.
-$\beta_{1i}$ and $\beta_{2i}$ are observer gains.
+The parameters $\beta_{1i}$ and $\beta_{2i}$ are observer gains.
 The estimation error dynamic is derived from (23) and (24):
@@
 where $\vartheta_{i}=\begin{bmatrix}x_{2i}-\hat{x}_{2i}\\ x_{3i}-\hat{x}_{3i}\\ \end{bmatrix}$, $A_{i}=\begin{bmatrix}-\beta_{1i}&1\\ -\beta_{2i}&0\\ \end{bmatrix}$, $E_{i}=\begin{bmatrix}0\\ 1\end{bmatrix}$.
-The observer gains are chosen such that the eigenvalues of the matrix $A_{i}$ are placed at $\omega_{o_{i}}$, where $\omega_{o_{i}}$ is the only tunable value known as the observer bandwidth [30].
+The observer gains are chosen such that the eigenvalues of the matrix $A_{i}$ are placed at $\omega_{o_{i}}$, where $\omega_{o_{i}}$ is the only tunable value known as the observer bandwidth in the ESO design [gao2003scaling].
 IV-C Controller Design
-$\hat{f}$ can be used to develop a disturbance rejection controller that complements the control law (13):
+The estimated disturbance $\hat{f}$ can be used to develop a disturbance rejection controller that complements the control law (15):
 […]
-where $\dot{q}_{d_{0}}$ is the nominal control from (13), $\dot{q}_{d_{f}}$ rejects disturbance, and they are summed to get $\dot{q}_{d}$.
+where $\dot{q}_{d_{0}}$ is the nominal control derived from (15), $\dot{q}_{d_{f}}$ is designed to cancel the disturbance, and they are summed together to obtain the control $\dot{q}_{d}$.
+$\bar{M}$ and $\bar{K}_{d}$ require only nominal values, not the true values.
+In Sec.
+IV-F, we provide more details on how to estimate them.
 IV-D Controller Stability Analysis
-Assumption 3
-There exists a positive known constant $l_{f}$ such that for any $x\in\mathcal{X}$, $u\in\mathcal{U}$, and $t\geq 0$, the following inequality holds:
-$\left|\dfrac{\partial f_{i}(t,x,u)}{\partial t}\right|\leq l_{f}.$ (25)
-Assumption 3 implies that $f$ is locally Lipschitz continuous with respect to time, and $\dot{f}$ is bounded.
 Theorem 1
-Given the robotic system (3) with kinematic control input, the ESO (22) with appropriate observer bandwidth $\omega_{o_{i}}$ for each joint, and the nominal control law (24) with appropriate tuned $k_{p}$ and $k_{d}$, we have the following properties: (i): the closed-loop system of (3) is bounded stable if each $\omega_{o_{i}}$ is finite; (ii): the closed-loop system of (3) is asymptotically stable if each $\omega_{o_{i}}$ approaches infinity.
+Given the robotic system (3) with kinematic control input, the ESO (24) with appropriate observer bandwidth $\omega_{o_{i}}$ for each joint, and the nominal control law (26) with appropriate tuned $k_{p}$ and $k_{d}$, we have the following properties:
+  1. 1.
+    the closed-loop system of (3) is bounded stable if each $\omega_{o_{i}}$ is finite;
+  2. 2.
+    the closed-loop system of (3) is asymptotically stable if each $\omega_{o_{i}}$ approaches infinity.
 Proof 1
-Since $A_{i}$ is Hurwitz, we have the following decomposition [31, Chapter 1.3]:
+Since $A_{i}$ is Hurwitz, we have the following decomposition [perko2013differential, Chapter 1.3]:
 $e^{A_{i}t}=P_{i}e^{\Lambda_{i}t}P_{i}^{-1},$ (27)
 where $P_{i}$ is a matrix whose columns are the eigenvectors of $A_{i}$, and $\Lambda_{i}$} is a diagonal matrix with the corresponding eigenvalues in the diagonal elements.
-Therefore, there exists a constant $c_{1_{i}}$ such that [31, Chapter 1.9]
+Therefore, there exists a constant $c_{1_{i}}$ such that [perko2013differential, Chapter 1.9]
 $\left\lVert e^{A_{i}t}\right\rVert\leq\left\lVert P_{i}\right\rVert\left\lVert e^{\Lambda_{i}t}\right\rVert\left\lVert P_{i}^{-1}\right\rVert=c_{1_{i}}\left\lVert e^{\Lambda_{i}t}\right\rVert\leq c_{1_{i}}\left\lVert e^{a_{i}t}\right\rVert,$ (28)
@@
 where $\epsilon=\begin{bmatrix}q_{i}-q_{i}^{\star}\\ \dot{q}_{i}-\dot{q}_{i}^{\star}\\ \end{bmatrix}$, $H=\begin{bmatrix}0&1\\ -k_{p_{i}}&-k_{d_{i}}\\ \end{bmatrix}$, and $\xi=\begin{bmatrix}0\\ \tilde{f}_{i}\\ \end{bmatrix}$.
-The matrix $H$ can be designed to be Hurwitz. Choose $V=\frac{1}{2}\epsilon^{T}Q\epsilon$ as the Lyapunov function candidate. Its derivative along (32) is given by:
-$\dot{V}=\tfrac{1}{2}(\dot{\epsilon}^{T}Q\epsilon+\epsilon^{T}Q\dot{\epsilon})=-\|\epsilon\|^{2}+\xi^{T}Q\epsilon.$ (33)
+The matrix $H$ can be designed to be Hurwitz, meaning there exists a symmetric positive definite matrix $Q$ such that
+$H^{T}Q+QH=-2I.$ (34)
+Choose $V=\frac{1}{2}\epsilon^{T}Q\epsilon$ as the Lyapunov function candidate. Its derivative along (33) is given by:
+$\begin{array}[]{ll}\dot{V}&=\frac{1}{2}(\dot{\epsilon}^{T}Q\epsilon+{\epsilon}^{T}Q\dot{\epsilon})\\ &=-\left\lVert\epsilon\right\rVert^{2}+\xi^{T}Q\epsilon.\end{array}$ (35)
 By applying the Cauchy–Schwarz inequality, we obtain:
@@
 IV-E Robust CBF for Safe Control
-By adding the disturbance term, system (11) becomes:
+By adding the disturbance term, system (13) can be written as follows:
 $\dot{x}=\psi(x)+g(x)u+\bar{g}f,$ (41)
@@
 Note that $u$ and $f$ are in the same channel, i.e., the $n\times n$ matrix of the upper part of $g(x)$ is a zero matrix and each row in the lower part of $g(x)$ includes an independent input, so $f$ is a matched disturbance vector.
-To define safety sets for system (39), we consider an $r$-times continuously differentiable function $h(x)$, where the relative degrees of $h(x)$ with respect to both $u$ and $f$ are $r$, given that $f$ is matched.
-A series of functions are defined as follows:
+To define safety sets for system (41), we consider an $r$-times continuously differentiable function $h(x)$, where the relative degrees of $h(x)$ with respect to both $u$ and $f$ are assumed to be $r$, given that $f$ is matched.
+A series of functions can be defined as follows:
 $\begin{array}[]{ll}h_{0}(x)=&h(x),\\ h_{1}(x)=&\dot{h}_{0}(x)+\gamma_{1}h_{0}(x),\\ &\vdots\\ h_{r}(x)=&\dot{h}_{r-1}(x)+\gamma_{r}h_{r-1}(x),\end{array}$ (42)
@@
 Then $\dot{h}_{r-1}(x)\geq 0$ for any $x\in\partial\mathcal{C}_{r-1}$.
-According to Nagumo’s theorem [32], since $x(0)\in\mathcal{C}_{r-1}$, we have $h_{r-1}(x)\geq 0$, $\forall t>0$.
+According to Nagumo’s theorem [blanchini2008set], since $x(0)\in\mathcal{C}_{r-1}$, we have $h_{r-1}(x)\geq 0$, $\forall t>0$.
 Iteratively, $\dot{h}_{0}(x)+\gamma_{1}h_{0}(x)\geq 0$, and we have $h_{0}(x)\geq 0$, $\forall t>0$, as $x(0)\in\mathcal{C}_{0}$.
@@
 Since the disturbance estimation error $\hat{f}-f$ is unknown in (47), the following assumption is made for the disturbance estimation error bound of ESO.
-The estimation error of $\hat{f}_{i}$ is as follows [29]:
+Assumption 3
+There exists a positive known constant $l_{f}$ such that for any $x\in\mathcal{X}$, $u\in\mathcal{U}$, and $t\geq 0$, the following inequality holds:
+$\left|\dfrac{\partial f_{i}(t,x,u)}{\partial t}\right|\leq l_{f}.$ (48)
+Assumption 3 implies that $f$ is locally Lipschitz continuous with respect to time.
+The estimation error of $\hat{f}_{i}$ is as follows [chen2022relationship]:
 $f_{i}(k)-\hat{f}_{i}(k)=p(k)*\Delta f_{i}(k),$ (49)
-where * represents convolution and
+where * represents convolution and [^2] For simplicity, let $\prod_{j=0}^{-1}(k+j)=1$.
 $\footnotesize\begin{array}[]{l}p(k)=\\ \begin{cases}1&1\leq k\leq r_{i}+1\\ \sum\limits_{\iota=1}^{r_{i}+1}\dfrac{1}{(\iota-1)!}\left(\prod\limits_{j=-\iota+1}^{-1}(k+j)\right)(1-\omega_{o_{i}})^{\iota-1}\omega_{o_{i}}^{k-\iota}&k\geq r_{i}+2.\end{cases}\end{array}$ (50)
-For simplicity, let $\prod_{j=0}^{-1}(k+j)=1$.
 Note that (49) is the disturbance estimation error formulated in the discrete-time domain.
 $f_{i}(k)=f_{i}(kT_{s})$, $\hat{f}_{i}(k)=\hat{f}_{i}(kT_{s})$, $\Delta f_{i}(k)=f_{i}((k+1)T_{s})-f_{i}(kT_{s})$, $T_{s}$ is the sample time used in calculating the error bound, and $r_{i}$ and $\omega_{o_{i}}$ are the relative degree of $\dot{q}_{i}$ (output measurement) with respect to $f_{i}$ (disturbance as input) in the $i$-th joint subsystem and the poles of the $i$-th ESO in the discrete-time domain, respectively.
-Readers are referred to our previous work [29] for a complete proof of this error bound.
+Readers are referred to our previous work [chen2022relationship] for a complete proof of this error bound, while a brief proof demonstrating that the sum of the series in (50) is bounded is provided in Appendix A.
 From the definition of derivative, Assumption 3, and (49), the disturbance estimation error bound of ESO is
 $|f_{i}(k)-\hat{f}_{i}(k)|\leq\Gamma_{i}(\omega_{o_{i}},T_{s})=\left(\sum\limits_{k=1}^{\infty}p(k)\right)l_{f}T_{s}.$ (51)
-As mentioned in [25], although (48) is derived in the discrete-time domain, the disturbance estimation error bound obtained from it is still the same as that from the continuous-time domain.
+As mentioned in [chen2023robust], although (51) is derived in the discrete-time domain, the disturbance estimation error bound obtained from it is still the same as that from the continuous-time domain.
 Since we need to utilize the disturbance estimation error bound in the continuous-time domain, $T_{s}$ should be small, such as 0.1 ms, in our experiment setting.
@@
 From Theorem 2, we need to prove $h_{r}(x)\geq 0$, $\forall t>0$. Substituting (46) and (51) into (45) yields
-$\begin{array}[]{r@{}l}h_{r}(x)=&L_{\psi}^{r}h(x)+L_{g}L_{\psi}^{r-1}h(x)u+L_{\bar{g}}L_{\psi}^{r-1}h(x)f\\ &+\sum\limits_{j=0}^{r-1}k_{j}h^{(j)}(x)\\ =&L_{\psi}^{r}h(x)+L_{g}L_{\psi}^{r-1}h(x)u+L_{\bar{g}}L_{\psi}^{r-1}h(x)\hat{f}\\ &-L_{\bar{g}}L_{\psi}^{r-1}h(x)(\hat{f}-f)+\sum\limits_{j=0}^{r-1}k_{j}h^{(j)}(x)\\ \geq&L_{\psi}^{r}h(x)+L_{g}L_{\psi}^{r-1}h(x)u+L_{\bar{g}}L_{\psi}^{r-1}h(x)\hat{f}\\ &-|L_{\bar{g}}L_{\psi}^{r-1}h(x)|\Gamma(\omega_{o},T_{s})+\sum\limits_{j=0}^{r-1}k_{j}h^{(j)}(x)\geq 0.\end{array}$ (50)
+$\begin{array}[]{r@{}l}h_{r}(x)=&L_{\psi}^{r}h(x)+L_{g}L_{\psi}^{r-1}h(x)u+L_{\bar{g}}L_{\psi}^{r-1}h(x)f\\ &+\sum\limits_{j=0}^{r-1}k_{j}h^{(j)}(x)\\ =&L_{\psi}^{r}h(x)+L_{g}L_{\psi}^{r-1}h(x)u+L_{\bar{g}}L_{\psi}^{r-1}h(x)\hat{f}\\ &-L_{\bar{g}}L_{\psi}^{r-1}h(x)(\hat{f}-f)+\sum\limits_{j=0}^{r-1}k_{j}h^{(j)}(x)\\ \geq&L_{\psi}^{r}h(x)+L_{g}L_{\psi}^{r-1}h(x)u+L_{\bar{g}}L_{\psi}^{r-1}h(x)\hat{f}\\ &-|L_{\bar{g}}L_{\psi}^{r-1}h(x)|\Gamma(\omega_{o},T_{s})+\sum\limits_{j=0}^{r-1}k_{j}h^{(j)}(x)\\ \geq&0.\end{array}$ (53)
 Then, we have $h_{r}(x)\geq 0$, $\forall t>0$ as $x(0)\in\mathcal{C}_{r}$.
@@
 Remark 2
-The key difference between our theoretical results in this work and our preliminary version [25] is that we have extended the robust CBF’s safety guarantee from a relative degree of one to a general case of higher order CBF.
-The safety specification requires the end-effector to maintain a non-negative distance from a virtual wall.
-The safety function is defined as $h(x)=y-y_{0}$, where $y$ and $y_{0}$ denote the positions of the end-effector and the virtual wall, respectively.
-It can be verified that $h(x)$ has a relative degree of two [26]:
+The fundamental difference between our main theoretical results in this work and our preliminary version [chen2023robust] is that we have extended the robust CBF’s safety guarantee from a relative degree of one to a general case of higher order CBF.
+Consider the safety specification for our end-effector: the relative distance between the end-effector and a virtual wall must be non-negative.
+The safety function can be defined as $h(x)=y-y_{0}$, where $y$ and $y_{0}$ are the positions of the end effector and the virtual wall, respectively.
+Remark 3
+We use this distance measurement as an example for simplification. Users can easily apply other safety specifications, such as the ball-like ones, without any fundamental difference.
+It can be verified that $h(x)$ has a relative degree of two with respect to system (13) by continuously taking its derivatives. A series of functions can be defined as follows [nguyen2016exponential]:
 […]
@@
 $\begin{array}[]{ll}h_{2}=&2\gamma\dot{y}+\gamma^{2}(y-y_{0})+\frac{d}{dt}\left(\frac{\partial\zeta}{\partial q}\right)\dot{q}\\ &+\frac{\partial\zeta}{\partial q}\Bigl(-\bar{M}^{-1}(\bar{C}\dot{q}+\bar{G}-\bar{K}_{d}\dot{q}_{d})+f)\Bigr).\end{array}$ (55)
-To simultaneously achieve tracking and safe control, the nominal controller for $\dot{q}_{d}$, designed in (24), should be minimally intervened. Thus, the following QP-CBF is constructed:
+To simultaneously achieve tracking and safe control, the nominal controller for $\dot{q}_{d}$, designed in (26), should be minimally intervened.
+That is, the nominal control should only be overridden by the safety controller when the system is unsafe.
+Thus, the following QP-based controller is constructed:
 $\dot{q}_{d_{\text{safe}}}=$ $\operatorname*{arg\,min}_{u\in\mathcal{U}}\|u-\dot{q}_{d}\|^{2}$ (56)
@@
 $+L_{\bar{g}}L_{\psi}h(x)\hat{f}-{|L_{\bar{g}}L_{\psi}h(x)|}\Gamma(\omega_{o},T_{s})$
-$+2\gamma\dot{h}(x)+\gamma^{2}h(x)\geq 0.$
-V Hardware Experiments
-We use the PUMA 500 robot for hardware experiments, as shown in Fig.
-3.
-It has three main joints and a spherical wrist, providing six degrees of freedom.
-$q_{1}$, $q_{2}$, and $q_{3}$ correspond to the waist, shoulder, and elbow joints, while $q_{4}$, $q_{5}$, and $q_{6}$ represent the wrist rotation, bend, and flange angles.
-Since our focus is task-space tracking (i.e., wrist center position), only the first three joints are used.
-Each joint is driven by a DC motor powered by an Adept Technology MV-19 Power Chassis.
-The control program is built in MATLAB Simulink, which generates C code and an SDF file, and runs via a dSPACE 1103 for control and data acquisition.
-Figure 3: PUMA 500 robot used as a testbed in our research.
-The inner-loop controller is configured as $\tau=-K_{p}({q}-{q}_{d})-K_{d}(\dot{q}-\dot{q}_{d})$, where $K_{p}=\mathrm{diag}([18,144,9])$ and $K_{d}=\mathrm{diag}([18,12,9])$.
-The gains are non-round values due to the actuator driver’s voltage-to-torque conversion ratio.
-The structure and true parameters remain unknown to our outer-loop controller.
+$+2\gamma\dot{h}(x)+\gamma^{2}h(x)\geq 0,$
+IV-F Determination of Nominal Control Gain
+We use the nominal model (14) throughout the experiments.
+A critical parameter in our controller design is the control gain in (26), $b_{0}\triangleq\bar{M}^{-1}\bar{K}_{d}$, which depends on the nominal inertia matrix $\bar{M}$ and the proportional gain of the inner-loop controller.
+We outline two possible approaches below for the determination of the (nominal) control gain.
+[…]
+Inertia matrix-based control gain tuning: Many common robotic manipulators, such as the PUMA robot used in our work, have available Universal Robot Description Format (URDF) files.
+A robot dynamics library can be used to parse the URDF and compute the inertia, Coriolis, and gravity matrices [carpentier2019pinocchio].
+We recommend that practitioners adopt this approach to obtain the nominal inertia matrix $\bar{M}$ for the control gain $b_{0}$ without requiring system identification.
+Once $\bar{M}$ is determined, tuning the control gain becomes solely a matter of adjusting $\bar{K}_{d}$.
+A practical rule of thumb for fine-tuning $\bar{K}_{d}$ is to start with a sufficiently large value, then gradually decrease it to improve tracking performance.
+A large gain generally leads to a larger but bounded tracking error while maintaining stability, whereas a small gain can reduce tracking error but increases the risk of instability if too small.
+The study in [xue2015performance] provides theoretical support for this trade-off.
+A useful tip for practitioners to simplify the fine-tuning process is to tune each joint individually rather than adjusting all coupled joints simultaneously.
+In our hardware experiments, where the control input is a lower-level voltage rather than direct torque control, the control gain becomes $\bar{M}^{-1}\bar{K}_{d}B$, where $B$ represents the mapping from voltage to torque.
+The inertia matrix $\bar{M}$ is derived from the same parameters used in our prior work [khalaf2019trajectory], though it can also be obtained from the URDF as mentioned earlier.
+Even if $K_{d}$ is assumed to be known in the experiments, $B$ still requires fine-tuning.
+Thus, control gain tuning remains necessary in the hardware experiments.
+Note that in the hardware experiments, the ground-truth control gain cannot be precisely determined due to the uncertainty in $\bar{M}$ (caused by estimation errors, payload variations, etc.) and the uncertainty in $B$ (as it is a simplified linear approximation between torque and voltage).
+We have conducted extensive experiments to demonstrate that with a broad range of $\bar{K}_{d}$—from 0.6 to 10 times its true value—the tracking error remains within 0.2 radians, indicating that gain tuning is practical, owing to its large working range, see Sec.
+VI-A.
+V Simulation Results
+In this section, we validate our method for safety assurance in a simulation environment.
+The advantage of using such a fully controlled environment is that we can obtain ground truth values for all kinematic and dynamic parameters, allowing us to directly compare the safety constraints of ours with the true constraints based on the actual parameters.
+The PUMA 500 robot model, as shown in Fig.
+3, is used for both simulation and the hardware experiments.
+The robot has three primary joints and a spherical wrist, providing a total of six degrees of freedom.
+$q_{1}$ represents the waist rotation angle, $q_{2}$ the shoulder rotation angle, and $q_{3}$ the elbow rotation angle, while $q_{4}$, $q_{5}$, and $q_{6}$ correspond to the wrist rotation angle, wrist bend angle, and flange angle, respectively.
+Our focus is tracking control in task space (i.e., the position of the wrist center); therefore, we only use the first three joints, as joints 4, 5, and 6 are revolute joints that do not affect the task space position.
+Figure 3: PUMA 500 robot used as a testbed in our research. The figure is from our previous work [khalaf2019trajectory].
+The robot model and system parameters here are adopted from our previous work [khalaf2019trajectory].
+The inner-loop controller is configured with a PD form, the structure of which is unknown to our outer-loop controller.
+$\tau=-K_{p}({q}-{q}_{d})-K_{d}(\dot{q}-\dot{q}_{d}).$ (57)
+where $K_{p}=\mathrm{diag}([1,1,1])$ and $K_{d}=\mathrm{diag}([1,1,1])$.
+The reference trajectory for joints $q_{1}$, $q_{2}$, and $q_{3}$ is $\begin{bmatrix}0.5\sin(t)&0.25\sin(2t)+0.5&0.25\sin(2t)+\frac{\pi}{2}\end{bmatrix}^{T}$.
+In the simulation, we demonstrate how ESO compensates for disturbances to ensure safety.
+A virtual wall is set at $y_{0}=\begin{bmatrix}y_{1}&-0.1&y_{2}\end{bmatrix}^{T}$, allowing free movement of the end-effector along the $x$ and $z$ axes, but constraining movement along the $y$ axis to a minimum value of 0.1m-0.1m.
+We follow the equation in IV-E for the CBF formulation, with the parameter $\gamma=10$ in the CBF.
+$\omega_{o}$ is set to $80$ rad/s for each joint.
+System (3) with ground truth parameters is used for simulation, neglecting the friction term, $F_{r}$ and the external disturbance, $F_{ext}$.
+Our nominal model is based on (14).
+The performance comparison is conducted using different CBFs: “True Model” refers to the CBF with the ground-truth model, “Nominal Model” refers to the CBF with the nominal model, and “Nominal Model with ESO” refers to the CBF with the nominal model plus disturbance compensation.
+Figure 4: The comparison of control bounds projected into each state with the true model (blue), nominal model (red), and nominal model with ESO (yellow).
+ESO can narrow the gap between the nominal model and the true model.
+Recall the CBF constraint in (44), which can be expressed as $u\geq LB(x)$ or $u\leq UB(x)$, where $LB(\cdot)$ and $UB(\cdot)$ represent the lower and upper bounds of the control input $u$.
+The determination of these bounds depends on the coefficient in front of $u$.
+In our case, the control input $u$ is $\dot{q}_{d}$, and the state $x$ includes the joint positions $q_{1}$, $q_{2}$, and $q_{3}$ and their derivatives.
+Additionally, note that the dimension of our control action is three.
+Fig.
+4 illustrates the bounds (lower or upper) of $\dot{q}_{d}$ based on the CBF under three scenarios: with ground truth disturbance $f$ (bounds shown as a solid blue line), robust CBF (ours, bounds shown as a dashed red line), and with only the nominal model (bounds shown as a yellow line).
+The y-axis represents the lower/upper bound values projected onto each state, and the x-axis represents time.
+As shown in the figure, the bounds generated by our approach closely match the true bounds, whereas the conventional CBF without robustness features fails to do so.
+The bounds play a critical role in ensuring safety for uncertain nonlinear dynamics.
+An inaccurate control bound in the CBF can result in overly conservative behavior or failure to ensure safety.
+In the hardware experiments, we will present more results with and without disturbance compensation.
+VI Hardware Experiments
+This section demonstrates the practical implementation of the proposed framework using the PUMA 500.
+Each joint of the PUMA 500 is actuated by a DC motor powered by an Adept Technology MV-19 Power Chassis.
+The control program is developed in MATLAB Simulink, which compiles the model and generates C code and an SDF file.
+Robot control and data acquisition are managed by a dSPACE 1103.
+Our robotic testbed used torque-mode servo drives, meaning that motor torques are given by $\tau=Bv$, where $B$ is a diagonal gain matrix and $v$ is an analog voltage command generated by the dSPACE system.
+Accordingly, the system model in (1) is modified to:
+${M}\ddot{q}+{C}\dot{q}+{G}+F_{r}=Bv+{J}^{T}F_{ext}.$ (58)
+The inner-loop controller is configured as
+$v=-K_{p}({q}-{q}_{d})-K_{d}(\dot{q}-\dot{q}_{d}).$ (59)
+where $K_{p}=\mathrm{diag}([1,12,1])$ and $K_{d}=\mathrm{diag}([1,1,1])$.
+Again, this structure is unknown to our outer-loop controller.
 The nominal second-order system model (14) is used for the controller.
@@
 The tracking performance of the nominal inner-loop controller is illustrated in Fig.
-4.
+5.
 Specifically, Fig.
-4a depicts the trajectory tracking performance in the joint space, while Fig.
-4b shows the trajectory tracking performance in the Cartesian space.
+5a depicts the trajectory tracking performance in the joint space, while Fig.
+5b shows the trajectory tracking performance in the Cartesian space.
 The results indicate that the inner-loop controller exhibits poor tracking performance.
 We will demonstrate the improvement achieved by adding our approach to the tracking and safe control tasks.
-We use the nominal model (12) throughout the experiments.
-A key parameter in the controller design is the control gain in (24), $b_{0}\triangleq\bar{M}^{-1}\bar{K}_{d}$, which depends on the nominal inertia matrix $\bar{M}$ and the inner-loop proportional gain.
-The nominal inertia $\bar{M}$ is adopted from [33].
-Following a common ESO tuning rule, we tune each joint individually by initially setting $b_{0}$ large to ensure stability and then gradually reducing it to improve tracking accuracy.
-The final nominal control gain is $\mathrm{diag}([20,40,10])$.
+The robot model and system parameters used here are adopted from our previous work [khalaf2019trajectory].
+The nominal $\bar{K_{d}}$ is set to $\mathrm{diag}([1,1,1])$, and the nominal $\bar{B}$ is estimated through open-loop testing and fine-tuned as $\mathrm{diag}([20,40,10])$.
 The observer and control bandwidths are set to $80~\mathrm{rad/s}$ and $10~\mathrm{rad/s}$, respectively, for each joint.
-We refer readers to [34] for additional instructions and theoretical support on tuning nominal control gain.
 (a) Joint space performance
@@
 The inner-loop controller can not track the reference trajectory.
-V-A Tracking Control
-We compare the tracking performance of ours, with [2].
-The reason is that we share the same minimal assumptions about the inner-loop controller, which, to the best of our knowledge, are the most general in the literature.
-V-A1 Scenario 1: Normal Case without Payload
-The robot tracks a reference trajectory without a payload.
-The tracking performance of the learning-based adaptive controller [2] and ours is shown in Fig.
-5.
+VI-A Robustness Testing under Different Nominal Gains
+Although the default nominal $\bar{K}_{d}$ is set equal to the true value, we have conducted extensive tests with different $\bar{K}_{d}$ values to evaluate the robustness of our approach when the nominal value is inaccurate.
 Fig.
-5a compares the reference (dashed) and actual joint trajectories of $q_{1}$, $q_{2}$, and $q_{3}$, where blue and red lines represent our method and [2], respectively.
-The x-axis denotes time (s), and the y-axis denotes joint angles (rad).
+6 shows the tracking error sensitivity to $\bar{K}_{d}$, while keeping other parameters fixed.
+The results demonstrate remarkable gain robustness, with the tracking error remaining below 0.2 radians across a wide range of $0.6\times,10\times]$ the true $K_{d}$ value.
+Figure 6: Comparison of tracking errors with different $K_{d}$ scaling factors using the nominal model (14).
+All results show a tracking error of less than 0.2 radians across a wide range of scaling factors: $0.6\times,10\times]$ the true $K_{d}$.
+The determination strategy follows the tuning guidance discussed in Section IV-F.
+Initial stabilization is achieved by setting $b_{0}$ sufficiently high, followed by a gradual reduction to improve tracking accuracy, until instability is observed at low gain values (e.g., unstable performance at $0.5K_{d}$ observed in the experiments).
+VI-B Tracking Control
+In this section, we compare the tracking performance of our approach, using the control law designed in (26) with that of the state-of-the-art learning-based adaptive control subject to closed architecture [ahanda2022adaptive].
+The reason for selecting [ahanda2022adaptive] for comparison is that we share the same minimal assumptions about the inner-loop controller, which, to the best of our knowledge, are the most general in the literature.
+In contrast, other works, such as [wang2019dynamic], assume more specific structures.
+We demonstrate the comparisons in a normal scenario without a payload and another scenario with different payloads.
+VI-B1 Scenario 1: Normal Case without Payload
+In this scenario, the robot is required to follow the prescribed reference trajectory without a payload.
+The tracking performance of both the learning-based adaptive control [ahanda2022adaptive] and ours is shown in Fig.
+7.
+The root mean squared errors (RMSEs) of the approaches under comparison are reported in Table II.
 Fig.
-5b shows the Cartesian trajectories in $x$, $y$, and $z$.
-The close alignment of the final and reference trajectories indicates that both controllers achieve accurate tracking despite model uncertainties and an imperfect inner loop.
-The proposed method achieves tracking performance comparable to the learning-based adaptive controller in the no-payload experiment.
+7a shows the joint reference and joint positions of three joints, $q_{1}$, $q_{2}$, and $q_{3}$.
+The dashed line is the reference trajectory, the blue line is the joint trajectory of ours, and the red line is the trajectory of [ahanda2022adaptive].
+The x-axis represents time (in seconds), and the y-axis denotes the angular position of the joints (in radians).
+Fig.
+7b illustrates the trajectory in the Cartesian space, with labeled axes $x$, $y$, and $z$.
+The close alignment of the final and reference trajectories suggests that both controllers can effectively track the desired trajectory despite model uncertainties and an imperfect inner-loop controller.
+The tracking performance of ours is comparable to that of the learning-based adaptive controller in this experiment without a payload.
 As shown in Fig.
-6, both approaches yield similarly smooth control signals.
-Transient responses occur initially, particularly in the second joint (Fig.
-5a), due to the offset between the initial and reference positions.
-The learning-based controller exhibits a longer transient phase (see Figs.
-5a and 6).
-This reflects a typical trade-off in adaptive control between transient duration and tracking accuracy.
+8, the control signal smoothness of the two approaches is also similar, with no significant difference.
+Transient states are observed at the beginning for both methods, especially for the second joint in Fig.
+7a, due to the distance between the initial robot joint positions and the reference.
+The learning-based adaptive controller takes longer in the transient phase (see the state transition in Fig.
+7a and the control variance in Fig.
+8).
+We notice that the adaptive method always involves a trade-off between transient time and tracking accuracy; i.e., to achieve better tracking performance, it requires a longer transient time, and vice versa.
 (a) Joint space comparison
@@
 Figure 7: Comparative tracking precision analysis: (a) Root Mean squared error (RMSE) in joint space: Proposed ($4.1\times 10^{-3}$, $2.76\times 10^{-2}$, $7.6\times 10^{-3}$) radians vs.
-[2] ($4.3\times 10^{-3}$, $2.7\times 10^{-2}$, $5.7\times 10^{-3}$) radians for each joint.
+[ahanda2022adaptive] ($4.3\times 10^{-3}$, $2.7\times 10^{-2}$, $5.7\times 10^{-3}$) radians for each joint.
 (b) Cartesian tracking error RMSE: Proposed ($2.3$ mm, $2.4$ mm, $15.5$ mm) vs.
-[2] ($2.9$ mm, $2.1$ mm, $16.1$ mm) for each joint.
+[ahanda2022adaptive] ($2.9$ mm, $2.1$ mm, $16.1$ mm) for each joint.
 Data sampled at 1 kHz over a 10 s trajectory execution.
+TABLE II: Comparison of RMSE in joint and Cartesian space
+Joint 1 Joint 2 Joint 3
+Root Mean Squared Error (RMSE) in Joint Space [rad]
+Proposed $4.1\times 10^{-3}$ $2.76\times 10^{-2}$ $7.6\times 10^{-3}$
+Proposed (zero-gravity) $1.3\times 10^{-3}$ $7.5\times 10^{-3}$ $2.4\times 10^{-3}$
+Adaptive [ahanda2022adaptive] $4.3\times 10^{-3}$ $2.7\times 10^{-2}$ $5.7\times 10^{-3}$
+x y z
+Cartesian Tracking Error RMSE [mm]
+Proposed $2.3$ $2.4$ $15.5$
+Proposed (zero-gravity) $1.5$ $0.7$ $4.0$
+Adaptive [ahanda2022adaptive] $2.9$ $2.1$ $16.1$
 Figure 8: Inner-loop control command comparison for $10$ s. $u_{1}$, $u_{2}$, and $u_{3}$ are the control inputs (voltages) for joint $q_{1}$, $q_{2}$, and $q_{3}$, respectively.
-V-A2 Scenario 2: Robustness Testing with Different Payloads
-We evaluate controller robustness with payloads of 1 kg, 1.5 kg, and 2 kg attached to the end-effector.
-Each payload experiment is repeated five times for ANOVA analysis, with identical controller and observer parameters.
+Comparison with and without gravity compensation: As noted in Assumption1, the gravity compensation term is optional in the inner-loop controller, since in practice, robotic arms may operate in either zero-gravity mode (where gravity is already compensated by the inner-loop controller) or non-zero-gravity mode.
+Without loss of generality, we choose the non-zero-gravity mode setting; see (57) and (59) for the simulation and hardware experiments, respectively.
+We also evaluate a variant of our method with a zero-gravity inner-loop controller (referred to as “Proposed (zero-gravity))” in Table II).
+The results show that both controllers achieve superior performance, demonstrating the robustness and generality of our approach across two different configurations.
+The zero-gravity mode yields slightly better results, which is reasonable because zero-gravity typically implies less disturbance to compensate for, making the control task easier.
+VI-B2 Scenario 2: Robustness Testing with Different Payloads
+In this scenario, we test the robustness of different controllers with payloads of 1 kg, 1.5 kg, and 2 kg attached to the end-effector, shown in Fig.
+9.
+Each experiment with a specific payload has been repeated five times for the ANOVA analysis.
+All controller and observer parameters have been kept the same.
 Fig.
-7a and Fig.
-7b present the tracking error ANOVA results for both methods.
-As shown earlier, our approach achieves a shorter transient time than the learning-based adaptive controller.
-For clarity, transient data (first 5 s) are excluded, and the plots show steady-state tracking errors (after 5 s).
+10a and Fig.
+10b show the ANOVA results of the tracking errors for the two approaches.
+As previously shown, our method exhibits a shorter transient time compared to the learning-based adaptive controller.
+For a more comprehensive comparison, the transient-phase data (first 5 seconds) were excluded, and the plots specifically present the steady-state tracking errors (from 5 seconds onward).
+Figure 9: Experimental setup for robustness testing. Payloads of 1 kg (left), 1.5 kg (middle), and 2 kg (right) are attached to the end-effector.
 (a) ANOVA analysis for ours
-(b) ANOVA analysis for learning-based adaptive control [2]
-Figure 7: Steady-state tracking error comparison under different payloads.
-(a) ANOVA of ours.
-(b) ANOVA of the learning-based adaptive control [2].
-The proposed method shows consistent tracking across payloads, as indicated by uniform box plots, while the adaptive method exhibits larger error variations, implying reduced robustness.
-Significance levels: ns ($p>0.05$), * ($p\leq 0.05$), ** ($p\leq 0.01$), *** ($p\leq 0.001$).
-Each experiment is repeated five times.
+(b) ANOVA analysis for learning-based adaptive control [ahanda2022adaptive]
+Figure 10: Steady-state tracking error comparison with different payloads.
+(a) ANOVA analysis of the proposed method.
+(b) ANOVA analysis of the learning-based adaptive control [ahanda2022adaptive].
+The proposed method maintains consistent tracking performance across different payloads, as evidenced by the uniform shape of the box plots; in contrast, the adaptive method shows significant variations in its error distribution, indicating compromised robustness.
+Significance levels: ns ($p>0.05$, not significant), * ($p\leq 0.05$), ** ($p\leq 0.01$), *** ($p\leq 0.001$).
+The experiments have been repeated five times.
 As shown in Fig.
-7a, the ANOVA results indicate that our method maintains statistically consistent errors across different payloads ($p>0.05$), demonstrating strong robustness, whereas the adaptive method shows significant variations with payload changes ($p<0.05$).
-In summary, our approach offers several advantages over the learning-based adaptive control [2].
-As shown in Fig.
-5, it achieves comparable tracking performance without a payload, lower steady-state error, and stable performance under varying payloads (Fig.
-7).
-It also exhibits a shorter transient response, as seen in the trajectories and control signals during the first second (Figs.
-5 and 6).
-Moreover, our design is simpler than a neural network, requiring fewer tunable parameters and offering greater interpretability.
-Finally, the proposed disturbance compensation supports both tracking and safe control with state constraints, whereas [2] cannot handle constraints.
-V-B Safe Control
-It is worth noting that $\gamma$ in (53) critically influences the CBF’s behavior.
-A larger $\gamma$ drives the trajectory closer to the safety set boundary (more aggressive), while a smaller $\gamma$ keeps it farther away (more conservative).
-To achieve a safe yet efficient controller, we set $\gamma=10$ in all experiments.
-We have conducted experiments in three scenarios: (1) Safety and Robustness Testing with Different Payloads, (2) Safety Testing Subject to External Disturbance, and (3) Safety Testing Subject to Gravity Disturbance.
-The first scenario will be described in detail, whereas the second and third scenarios are demonstrated in our accompanying video.
-A virtual boundary is defined at $y_{0}=\begin{bmatrix}y_{1}&-0.1&y_{3}\end{bmatrix}^{T}$ as a spatial constraint for the end effector.
-It moves freely along the $x$ and $z$ axes but is constrained along the $y$ axis with a lower limit of 0.1-0.1 m.
-The system is tested with payloads of $1$, $1.5$, and $2$ kg attached to the end effector.
-Three approaches are compared: (1) conventional CBF with a nominal model (see Fig.
-8); (2) our robust CBF with a nominal model and an ESO for disturbance estimation, compared to the DOB-CBF [16] (see Fig.
-9a); and (3) our robust CBF with a nominal model, ESO for disturbance estimation, and estimation error bound, compared to the DOB-CBF with estimation error bound (see Fig.
-9b).
+10a, the significance testing from the ANOVA demonstrates that our method maintains statistically consistent error distributions across varying payload conditions ($p>0.05$), indicating strong robustness.
+In contrast, the adaptive methods exhibit significant variations due to payload changes ($p<0.05$).
+In summary, our approach demonstrates significant advantages compared to the learning-based adaptive control [ahanda2022adaptive]:
+  1. 1.
+    As shown in Table II, the tracking performance of ours is comparable to that of the learning-based adaptive controller in this experiment with- out a payload.
+  2. 2.
+    Our approach exhibits lower steady-state error and achieves stable and consistent tracking performance under different payloads, as shown by the ANOVA test in Fig. 10.
+  3. 3.
+    Our approach has a shorter transient time, as shown in the trajectories and control signals during the first second in Fig. 7 and Fig. 8, respectively.
+  4. 4.
+    Our design is much simpler than a neural network, with far fewer parameters to tune and better interpretability, making it easier for users to perform troubleshooting.
+  5. 5.
+    Our disturbance compensation is not only used for tracking but can also be applied in safe control design with state constraints, whereas [ahanda2022adaptive] cannot handle any constraints.
+However, we also acknowledge that [ahanda2022adaptive] unifies all components within a learning framework, whereas ours approach requires an additional step for estimating the nominal control gain.
+VI-C Safe Control
+In this section, we validate the safe control performance of our approach.
+It is essential to highlight that $\gamma$ in (56) is a crucial parameter related to the CBF’s behavior.
+A large $\gamma$ makes the trajectory approach closer to the safety set’s boundary, meaning a more aggressive behavior.
+A small $\gamma$ enforces the trajectory to stay far away from the safety set’s boundary, meaning a more conservative behavior.
+To design a safe and efficient controller that is desired in practice, we set $\gamma$ to a significantly high value of $10$ for all the experiments in this section.
+As previously mentioned, disturbances can either help push the trajectory back into the interior of the safety set or cause it to escape the safety set.
+Both of these behaviors are undesirable.
+We will showcase these three scenarios in the following experiments: (1) Safety and Robustness Testing with Different Payloads, (2) Safety Testing Subject to External Disturbance, and (3) Safety Testing Subject to Gravity Disturbance.
+VI-C1 Safety and Robustness Testing with Different Payloads
+A virtual boundary is set at $y_{0}=\begin{bmatrix}y_{1}&-0.1&y_{3}\end{bmatrix}^{T}$ as a spatial constraint for the end effector.
+The end effector is allowed unrestricted movement along the $x$ and $z$ axes but is constrained along the $y$ axis, with a lower limit set at 0.1-0.1 m.
+The system is evaluated with payloads of $1$ kg, $1.5$ kg, and $2$ kg attached to the end effector.
+We will present three approaches for comparison: (1) conventional CBF with a nominal model (see the safety function $h$ in Fig.
+11); (2) our robust CBF with a nominal model and an ESO for disturbance estimation, comparared to the DOB-CBF [dacs2022robust] (see Fig.
+12a); and (3) our robust CBF with a nominal model, ESO for disturbance estimation, and estimation error bound, compared to the DOB-CBF with estimation error bound (see Fig.
+12b).
 Figure 11: Safety boundary: $\bm{y_{0}}=\begin{bmatrix}\bm{y_{1}}&\bm{-0.1}&\bm{y_{3}}\end{bmatrix}^{T}$.
@@
 Fig.
-8 shows the safety performance of the CBF using only the nominal second-order model (12), which neglects model errors, inner-loop uncertainty, and external disturbances.
-The safety function $h$ varies under different payloads, revealing limited robustness.
-We implement the DOB-CBF [16] for comparison.
-The DOB-CBF estimates $b_{e}(x,d)=\frac{\partial\zeta}{\partial q}f$ in the CBF, while our method directly estimates $f$ in the original dynamics.
-Thus, the two observers target different quantities, and their gains are not directly comparable.
-For fairness, excluding error bounds, we tune both observers’ gains to achieve similar minimum $h$ magnitudes, indicating comparable safety performance.
+11 illustrates the safety performance of CBF with only a nominal second-order system model (14).
+This nominal model does not account for any uncertainty arising from model errors, an uncertain inner-loop controller, or external disturbances.
+The safety function $h$ exhibits variations due to external disturbances from varying payloads, indicating a lack of robust performance.
+The details of implementing the DOB-CBF [dacs2022robust] have been included in Appendix B.
+The DOB-CBF estimates $b_{e}(x,d)=\frac{\partial\zeta}{\partial q}f$ in CBF, whereas our approach directly estimates $f$ in the original dynamic system.
+This leads to a fundamental difference: the two observers estimate different quantities, making direct comparison difficult.
+In particular, the key parameters—observer gains—cannot be directly compared because they pertain to two fundamentally different observers.
+To ensure fairness, without including error bounds, we tune the observer gains for both approaches such that they yield similar minimum $h$ values (in terms of magnitude), suggesting comparable safety performance without error bounds.
 Fig.
-9a compares the $h$ functions of the DOB-CBF and the proposed method, showing that:
+12a illustrates the $h$ function comparison between the DOB-CBF and the proposed method.
+It is observed that:
   1. 1.
@@
   2. 2.
-    A slight safety violation occurs due to disturbance estimation error, with minimum values of 0.597-0.597 mm for DOB-CBF and 0.366-0.366 mm for the proposed method, highlighting the need to account for estimation error bounds in both.
-We collect experimental data for $\hat{b}_{e}(x,d)$ and $\hat{f}$ from the DOB and ESO, respectively.
-Their finite differences yield the maximum rates of change, used as estimation error bounds.
+    Due to disturbance estimation error, a slight violation of the safety specification occurs, with the lowest point reaching 0.597-0.597 mm for DOB-CBF and 0.366-0.366 mm for the proposed method.
+This necessitates the consideration of estimation error bounds for both methods.
+We collect experimental data for the estimated $\hat{b}_{e}(x,d)$ and $\hat{f}$ from the DOB and ESO, respectively.
+Their finite differences are computed to obtain the maximum values, which serve as bounds on their rates of change.
 Fig.
-9b shows the resulting $h$ values after incorporating these bounds.
-Compared with Fig.
-9a, we observe that:
+12b illustrates the resulting $h$ values after incorporating the estimation error bounds of the two approaches.
+Compared to the previous experiment shown in Fig.
+12a, we have the following observations:
   1. 1.
-    The trajectories remain safe and consistent across different payloads, demonstrating the robustness of both methods.
+    The trajectories are safe and consistent under different payloads, indicating robustness of both methods.
   2. 2.
-    Incorporating the estimation error bound raises $h$, keeping trajectories strictly within the safety set (positive $h$ values).
+    The incorporation of the estimation error bound elevates the value of $h$, making the trajectories always stay in the safety set; see the positive $h$ function values.
   3. 3.
-    With error bounds, our method stays closer to the safety boundary than [16], indicating lower conservatism.
-This may be because [16] estimates $b_{e}(x,d)=\frac{\partial\zeta}{\partial q}f$, which includes a state-dependent Jacobian, making it more sensitive to measurement noise and compounding estimation errors.
-VI Conclusion and future work
-This paper presents a unified framework that improves tracking performance and ensures safe kinematic control for closed-architecture robotic systems.
-The proposed method integrates seamlessly into existing cascade control structures as an add-on, enabling robust and safe control under uncertain dynamics, external disturbances, and unknown inner-loop controllers.
-Stability and safety are theoretically guaranteed, and real-time validation is conducted on the PUMA 500 industrial robot.
-Future work will extend the framework to visual servoing systems, addressing uncertainties in the vision pipeline.
+    When incorporating error bounds, our approach remains closer to the safety boundary than that of [dacs2022robust], indicating reduced conservatism.
+A possible explanation is that [dacs2022robust] estimates $b_{e}(x,d)=\frac{\partial\zeta}{\partial q}f$, which involves a state-dependent Jacobian matrix.
+In a real robot with measurement noise, this introduces additional uncertainty, on top of the uncertainty from the disturbance $f$, potentially resulting in compounded estimation error.
+However, we acknowledge that the difference is not large, typically within the order of millimeter in $h$.
+For practitioners using our approach, we offer the following recommendations regarding whether to include the error bound: First, if the CBF and disturbance compensation already yield satisfactory results—i.e., ensuring safety without being overly conservative—then incorporating the error bound may be unnecessary, particularly when the disturbance rate of change cannot be estimated within a tight range.
+Second, if performance is compromised due to estimation errors, then including the error bound is advised.
+However, it is crucial that the disturbance rate of change be carefully estimated; overly loose estimates can lead to over-conservatism.
+VI-C2 Safety Testing Subject to External Disturbance
+In this scenario, a constant 6V is applied to $q_{1}$, the waist joint, intentionally generating a continuous external perturbation that pushes the manipulator’s end-effector toward the virtual wall to evaluate the controller’s response.
+As shown in (58), the motor torque is controlled by an analog voltage.
+Figure 13: Safety boundary: $\bm{y_{0}}=\begin{bmatrix}\bm{y_{1}}&\bm{-0.1}&\bm{y_{3}}\end{bmatrix}^{T}$.
+Trajectory comparison of CBF subject to external disturbance.
+The trajectory (blue) always stays in the safe zone.
+The unsafe trajectory (red) crosses the safety boundary.
+Fig.
+13 illustrates a trajectory comparison of CBFs with and without disturbance compensation.
+The red line shows the unsafe trajectory of the CBF that uses only the nominal model (14).
+While the tracking performance is satisfactory due to the nominal controller’s disturbance rejection capacity, the CBF is expected to intervene when the reference trajectory crosses the safety boundary.
+Unfortunately, the final trajectory still crosses the safety boundary because the CBF does not account for the external disturbance pushing the end-effector toward the unsafe region.
+In contrast, as shown in the blue line, our approach ensures high-performance tracking when safety is not a concern (see area $y=[0,0.3]$) and safe control, as the trajectory never crosses the safety boundary.
+VI-C3 Safety Testing Subject to Gravity Disturbance
+In this scenario, we set the virtual boundary as $y_{0}=\begin{bmatrix}y_{1}&y_{2}&-0.01\end{bmatrix}^{T}$.
+The end effector is allowed unrestricted motion along the $x$- and $y$-axes, but movement along the $z$-axis is constrained by a lower limit of 0.01-0.01 m.
+The disturbance caused by gravity will inevitably push the end effector toward the virtual boundary, leading to unsafe behavior.
+Remember that our nominal model (14), which has minimal model information, treats gravity as a disturbance.
+Figure 14: Safety boundary: $\bm{y_{0}}=\begin{bmatrix}\bm{y_{1}}&\bm{y_{2}}&\bm{-0.01}\end{bmatrix}^{T}$.
+Trajectory comparison of CBFs subject to gravity disturbance.
+The safe trajectory (blue, ours) remains within the safe zone, while the unsafe trajectory (red, without disturbance compensation) crosses the safety boundary.
+Fig.
+14 shows the trajectory comparison of CBFs subject to gravity disturbance.
+The red line illustrates the unsafe trajectory of a CBF using only the nominal model (14).
+In contrast, our robust CBF significantly improves performance, as shown by the blue line, which consistently remains within the safe zone.
+Computation complexity: The computational complexity is not the reason for removing joints 4, 5, and 6.
+Their removal is solely due to their redundancy in the 3D position control of the end-effector (i.e., the position of the wrist center).
+We break down the essential computations in our framework: (i) The observer in (24) involves only the simple integration of ordinary differential equations.
+It can easily scale to higher dimensions.
+(ii) The disturbance rejection controller in (26) has an explicit form without heavy computation.
+(iii) Our robust CBF in (56) is still based on standard convex Quadratic Programming (QP) solving, which is scalable to higher dimensions.
+The only additional terms are from the disturbance computed from the observer and the disturbance error bound (51).
+The disturbance error bound is computed offline.
+We emphasize that our hardware implementation operates at a 1 kHz frequency for both measurement sampling and control in dSPACE.
+Possible extensions to other robotic systems: Our approach has the potential to be extended to other robotic systems, particularly commercial robots that face the same challenge addressed in this work: closed architectures with only kinematic controllers and inner-loop controllers implemented as unmodifiable firmware.
+The rationale behind this possible extension is that our dynamic system follows a general second-order Euler–Lagrange form, which is widely used in robotic motion control systems.
+In addition, the cascade control structure is common, with a kinematic controller in the outer loop and an inner-loop torque controller that typically includes a proportional term and an unknown function—potentially incorporating integral and dynamic compensation terms.
+However, we acknowledge that further customization may be necessary; for example, the presence of mismatched disturbances, which are common in many underactuated robots, poses additional challenges.
+Our recent work has demonstrated promising results in rejecting such mismatched disturbances in underactuated drones [chen2024quadrotor].
+VII Conclusion and future work
+A unified framework that enhances tracking performance and ensures safe kinematic control for robotic systems with closed architecture is proposed in this paper.
+Compared to the state-of-the-art method in [ahanda2022adaptive], which is used solely for tracking, our approach introduces safety features.
+The proposed approach can be easily integrated into the existing cascade control framework as an add-on, providing robust and safe control in the presence of various uncertainties, including uncertain dynamics, external disturbances, and an uncertain inner-loop controller.
+Theoretical results of stability analysis and provable safety guarantees are presented.
+The framework has been validated in real-time on the PUMA 500 industrial robot.
+In future work, we plan to extend our approach to a visual servoing control system to address additional challenges, particularly when uncertainty in the vision system needs to be considered.
+Appendix A Proof of the Bounded Series in (51)
+The sum of the infinite series in (51) can be expressed as
+$\begin{array}[]{r@{}l}\sum\limits_{k=1}^{\infty}p(k)=&r_{i}+1+\sum\limits_{\iota=1}^{r_{i}+1}\dfrac{1}{(\iota-1)!}(1-\omega_{o_{i}})^{\iota-1}\\ &\cdot\sum\limits_{k=1}^{\infty}\left(\prod\limits_{j=-\iota+1}^{-1}(k+j)\right)\omega_{o_{i}}^{k-\iota},\end{array}$ (60)
+where the finite sum on the right-hand side is clearly bounded.
+The key step is to show that the second infinite sum converges.
+Note that since $j<0$, $k>0$, $\iota\geq 1$ and $\omega_{o_{i}}>0$, we can write
+$\sum\limits_{k=1}^{\infty}\left(\prod\limits_{j=-\iota+1}^{-1}(k+j)\right)\omega_{o_{i}}^{k-\iota}\leq\sum\limits_{k=1}^{\infty}k^{\iota-1}\omega_{o_{i}}^{k-\iota}.$ (61)
+Applying the ratio test, we have
+$\lim_{k\to\infty}\frac{(k+1)^{\iota-1}\omega_{o_{i}}^{k+1-\iota}}{k^{\iota-1}\omega_{o_{i}}^{k-\iota}}=\lim_{k\to\infty}\left(\frac{k+1}{k}\right)^{\iota-1}\omega_{o_{i}}=\omega_{o_{i}}.$ (62)
+Since $\omega_{o_{i}}<1$ in the discrete-time domain, the infinite series $\sum\limits_{k=1}^{\infty}k^{\iota-1}\omega_{o_{i}}^{k-\iota}$ converges. This completes the proof.
+Appendix B Implementation Details of DOB-CBF [dacs2022robust] for Comparison
+The DOB proposed in [dacs2022robust] is designed for a system with relative degree one. In other words, it estimates the effect of disturbances on the dynamics of $h$, i.e.,
+$\underbrace{\ddot{h}(x)}_{\dot{h}^{{}^{\prime}}(x)}=\underbrace{L_{f}^{2}h(x)+L_{g}L_{f}h(x)u}_{a_{e}(x,u)}+\underbrace{L_{g}L_{f}h(x)f}_{b_{e}(x,f)},$ (63)
+where $a_{e}(x,u)$ is known and $b_{e}(x,f)$ needs to be estimated.
+In this paper, $\ddot{h}$ is expressed as follows:
+$\begin{array}[]{ll}\ddot{h}&=\frac{d}{dt}\left(\frac{\partial\zeta}{\partial q}\right)\dot{q}+\frac{\partial\zeta}{\partial q}\Bigl(-\bar{M}^{-1}(\bar{C}\dot{q}+\bar{G}-\bar{K}_{d}\dot{q}_{d})+f)\Bigr)\\ &=\underbrace{\frac{d}{dt}\left(\frac{\partial\zeta}{\partial q}\right)\dot{q}+\frac{\partial\zeta}{\partial q}\Bigl(-\bar{M}^{-1}(\bar{C}\dot{q}+\bar{G}-\bar{K}_{d}\dot{q}_{d})\Bigr)}_{a_{e}(x,u)}+\underbrace{\frac{\partial\zeta}{\partial q}f}_{b_{e}(x,f)}.\end{array}$ (64)
+Using the DOB-CBF approach, the unknown term to be estimated is $b_{e}(x,f)=\frac{\partial\zeta}{\partial q}f$. A DOB is designed as follows:
+$\begin{cases}\begin{aligned} \hat{b}_{e}&=k_{b}\dot{h}-\chi,\\ \dot{\chi}&=k_{b}(a_{e}(x,u)+\hat{b}_{e})\\ &=k_{b}(a_{e}(x,u)+k_{b}\dot{h}-\chi),\end{aligned}\end{cases}$ (65)
+where $k_{b}$ is the observer gain, $\dot{h}=\left(\frac{\partial\zeta}{\partial q}\right)\dot{q}$, and $\zeta$, $q$ and $\dot{q}$ are obtained from the measurements.
+Suppose $b_{e}(x,f)$ is differentiable with respect to $t$, with an upper bound on its rate of change, $|\dot{b}_{e}(x,f)|\leq b_{h}$.
+Then, the steady-state estimation error bound of $\hat{b}_{e}$ is given by $\frac{b_{h}}{k_{b}}$.